### v2

# SADA: Safe and Adaptive Aggregation of Multiple Black-Box Predictions in Semi-Supervised Learning

Abstract
Semi-supervised learning (SSL) arises in practice when labeled data are scarce or expensive to obtain, while large quantities of unlabeled data are readily available. With the growing adoption of machine learning techniques, it has become increasingly feasible to generate multiple predicted labels using a variety of models and algorithms, including deep learning, large language models, and generative AI. In this paper, we propose a novel approach that safely and adaptively aggregates multiple black-box predictions of uncertain quality for both inference and prediction tasks. Our method provides two key guarantees: (i) it never performs worse than using the labeled data alone, regardless of the quality of the predictions; and (ii) if any one of the predictions (without knowing which one) perfectly fits the ground truth, the algorithm adaptively exploits this to achieve either a faster convergence rate or the semiparametric efficiency bound. We demonstrate the effectiveness of the proposed algorithm through small-scale simulations and two real-data analyses with distinct scientific goals. A user-friendly R package, sada, is provided to facilitate practical implementation.
Key Words: Black-box models; Prediction-powered inference; Semiparametric efficiency; Semi-supervised learning.
1 Introduction
In many scientific and industrial applications, acquiring gold-standard labels can be expensive, time-consuming, or require specialized expertise, while collecting unlabeled data is relatively cheap and scalable. This disparity motivates semi-supervised learning (SSL), a paradigm that combines scarce labeled observations with large amounts of unlabeled samples (Zhu, 2005; van Engelen & Hoos, 2020). Over the past decades, a wide range of SSL algorithms have been developed to exploit abundant unlabeled data to improve model performance, for example by using a pre-trained model to assign pseudo labels to unlabeled samples (Lee, 2013; Zhu et al., 2023) or by propagating information through relationships between data points (Jebara et al., 2009; Subramanya & Talukdar, 2014).
With the rapid evolution of machine learning and generative AI, it has become much easier than ever to generate predicted labels using powerful tools such as large language models (LLMs). These predictions can often be obtained at substantially lower cost and with much shorter turnaround time than human labels. Consequently, leveraging machine-learned predictions in SSL has become an increasingly important practical demand. In the context of statistical inference, Angelopoulos et al. (2023) introduced prediction-powered inference (PPI), a framework that enables valid inference without imposing any assumptions on the predictions. However, when the predictive model is of low quality, PPI may have worse estimation efficiency than labeled-only baselines. This limitation has motivated further work aimed at improving the efficiency or integrating it with ideas from other statistical and machine learning frameworks (Angelopoulos et al., 2024; Zrnic & Candès, 2024; Fisch et al., 2024; Ji et al., 2025; Miao et al., 2025).
Despite these advances, several gaps and challenges remain when one considers modern SSL pipelines in which multiple predicted labels are available. First, current PPI methods are designed for a single predictive model, whereas in practice it is common to obtain predictions from multiple models or sources. Although one may heuristically select the seemingly best model or apply ensembling strategies, there is no general framework within SSL for aggregating multiple predictions in a principled manner. Second, outputs from different models, such as GPT, Llama, or DeepSeek, can differ substantially, and the quality of predictions from black box models can be highly variable. In particular, low-quality or poorly calibrated predictions can introduce significant noise, increasing variance and leading to unreliable inference. Third, most existing work targets statistical inference for the paramter of interest; however, in many applications such as image classification, practitioners are primarily concerned with predictive model accuracy rather than parameter estimation. Current PPI methods and their extensions do not directly address how to improve prediction accuracy while retaining rigorous guarantees.
In this paper, we introduce SADA, a novel and comprehensive procedure that safely and adaptively aggregates multiple predicted labels originating from diverse black-box models for both inference and prediction tasks. Our approach is characterized by two key features: safety and adaptivity. Safety ensures that the proposed estimator never performs worse than a labeled-only baseline, regardless of the quality of the predictions. Specifically, for inference, the SADA estimator guarantees an asymptotic variance no larger than that of the estimator using labeled data only, a property that is generally not ensured by PPI (Angelopoulos et al., 2023) or PPI++ (Angelopoulos et al., 2024). For prediction, the excess risk of the SADA estimator converges at a rate no slower than that of the labeled-only baseline. Adaptivity means that our algorithm automatically aggregates multiple predictions by assigning them appropriate weights, leading to more efficient inference or improved predictive accuracy. Notably, in cases where one prediction is perfectly accurate, SADA automatically identifies and leverages it to attain a faster convergence rate or the semiparametric efficiency bound.
Our contributions are threefold. First, we propose a principled and fully data-driven procedure for aggregating multiple predicted labels of uncertain quality in a semi-supervised learning setting, accompanied by theoretical guarantees, which greatly enhances and broadens the applicability of PPI-based methods in practice. Second, we derive the asymptotic representations and excess risk bounds for the SADA estimator, which, make explicit how the method operates, demonstrate how it protects against potentially low-quality predictions, and clarify how it automatically borrows strength from informative predictions without requiring prior knowledge or formal testing of the quality of each model. Third, we illustrate these properties through intuitive explanations, simulation studies, and two real-data applications with distinct tasks that demonstrate the practical usefulness of the approach. We also provide a user-friendly R package, sada, to facilitate implementation of the proposed algorithm.
1.1 Related work
Semi-supervised learning (SSL).
Over the past decades, a wide range of SSL algorithms has been proposed. These methods differ in the assumptions, in how they utilize unlabeled data, and in their relationship to supervised learning approaches (van Engelen & Hoos, 2020). Broadly, these algorithms can be categorized into two types: inductive and transductive methods. Inductive methods, similar to supervised learning, use a pre-trained model to assign labels to unlabeled data. Examples include self-training (Yarowsky, 1995; Lee, 2013; Berthelot et al., 2019, 2020), co-training (Blum & Mitchell, 1998; Wang & Zhou, 2010; Deng & Guo, 2011), pseudo-labeled boosting methods (Zhou, 2012), unsupervised preprocessing (Sheikhpour et al., 2017). In contrast, transductive methods do not produce a generalizable model; instead, they predict labels by directly propagating information through connections between data points. It typically defines a graph over all data points, both labeled and unlabeled, encoding the pairwise similarity of data points with possibly weighted edges (Jebara et al., 2009; Liu et al., 2012; Subramanya & Talukdar, 2014). Additionally, significant progress has been made in recent years to understand, as well as how to leverage, the statistical benefits of the unlabeled data. Chakrabortty & Cai (2018) and Azriel et al. (2022) studied linear regression problems within the SSL framework and proposed estimators that are more efficient than ordinary least squares (OLS) which relies solely on labeled data. Song et al. (2023) further extended this framework to general M-estimation problems. The methodology has also been adapted to high-dimensional settings, where the number of features exceeds the sample size (Zhang et al., 2019; Cai & Guo, 2020; Zhang & Bradic, 2022; Deng et al., 2024). Applications of SSL have expanded beyond statistical models to include both 2D computer vision tasks (Jeong et al., 2019; Liu et al., 2021; Zhou et al., 2022) and 3D object detection problems (Wang et al., 2021; Park et al., 2022; Liu et al., 2023).
Prediction-powered inference (PPI).
In the past few years, a growing body of research has focused on enhancing statistical inference by incorporating predictions from black-box models (Wang et al., 2020; Motwani & Witten, 2023). In particular, Angelopoulos et al. (2023) introduced prediction-powered inference (PPI), a framework that enables valid inference even when the predictive model is of low quality. However, PPI might be less efficient than the labeled-only baseline, motivating follow-up works aimed at improving its efficiency or combining it with other statistical and machine learning ideas. Examples include PPI++ (Angelopoulos et al., 2024), cross PPI (Zrnic & Candès, 2024), stratified PPI (Fisch et al., 2024), recalibrated PPI (Ji et al., 2025).
In related work, Zhu et al. (2023) proposed a doubly robust self-training method that achieves faster convergence rates when predictions are highly accurate. Miao et al. (2025) introduced a post-prediction adaptive inference approach that ensures valid statistical inference without relying on assumptions about the predictions. Gan et al. (2024) explored a broader class of imputed loss functions to enhance modeling flexibility and efficiency. Gronsbell et al. (2025) focused on inference under squared error loss, situating PPI within the broader context of semiparametric theory. Bartolomeis et al. (2025) introduced a framework that integrates the predictions from multiple foundation models with randomized experiments while preserving valid statistical inference.
Missing data and causal inference.
SSL is also closely related to missing data and causal inference (Rubin, 1974, 1976). When missingness is at random, a rich variety of methods has been conducted to handle this, such as likelihood-based inference (Ibrahim, 1990), imputation (Rubin, 2004), and semiparametric methods (Robins et al., 1994; Zhao et al., 1996). For more complex missing-not-at-random settings, identification and estimation methods are also well developed, such as outcome selection models (Heckman, 1979), graphical models (Fay, 1986; Ma et al., 2003), sensitivity analysis techniques (Robins et al., 2000), and approaches based on instrumental variables or shadow variables (Wang et al., 2014; Zhao & Shao, 2015; Zhao & Ma, 2022; Miao et al., 2024).
1.2 Paper organization
The rest of this paper is organized as follows. Section 2 introduces the problem setup and data structure. Section 3 presents the SADA approach for inference and develops its theoretical properties, while Section 4 adapts the procedure to the prediction task. Section 5 constructs a small-scale simulation study to evaluate the finite-sample performance of the proposed methods. Section 6 applies the proposed methods to two real-data analyses with distinct tasks. Section 7 concludes the paper. All proofs are provided in the Supplementary Material.
1.3 Notations
The following notations are used throughout the paper. All vectors are assumed to be column vectors unless otherwise specified. Let $\mathbf{u}^{\otimes 2}=\mathbf{u}\mathbf{u}^{\mathrm{\scriptscriptstyle T}}$ for a vector $\mathbf{u}$. We denote $\mathbf{A}\preceq\mathbf{B}$ for two symmetric square matrices $\mathbf{A}$ and $\mathbf{B}$ if $\mathbf{B}-\mathbf{A}$ is positive semi-definite. We generally use capital letters to denote random variables, calligraphic letters to denote their support, and lowercase letters to denote their realizations. We use $\mathbb{P}$ to denote the probability measure and $\mathbb{E}$ to denote the expectation. For two random vectors $\mathbf{U}$ and $\mathbf{V}$, let $\operatorname{cov}(\mathbf{U},\mathbf{V})=\mathbb{E}[\{\mathbf{U}-\mathbb{E}(\mathbf{U})\}\{\mathbf{V}-\mathbb{E}(\mathbf{V})\}^{\mathrm{\scriptscriptstyle T}}]$ and $\operatorname{var}(\mathbf{U})=\operatorname{cov}(\mathbf{U},\mathbf{U})$.
2 Setup
In this paper, we consider two types of scientific goals: inference and prediction. Let $Y$ be a scaler outcome and $\mathbf{X}$ be a vector of covariates. For the inference problem, the target is to estimate and draw inference on a $p$-dimensional parameter, ${\bm{\theta}}^{*}\in\Theta\subset\mathbb{R}^{p}$, defined through an estimating equation
$\mathbb{E}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})\}=\mathbf{0},$ (1)
where $\mathbf{s}(\mathbf{x},y;{\bm{\theta}})$ is a user-specified $p\times 1$ vector-valued function. This formulation is broad and include many commonly studied parameters, such as outcome means, quantiles, least squares coefficients, and other functionals of $\mathbf{X}$ and $Y$, including the unique minimizer of a loss function or the maximizer of a criterion function (van der Vaart, 1998). The goal is to construct an estimator of $\bm{\theta}$$^{*}$ that is as close as possible to the true value $\bm{\theta}$$^{*}$.
By contrast, a prediction problem is defined by a hypothesis family $\{f_{\bm{\theta}}:\mathcal{X}\to\mathcal{Y}\}$ parameterized by ${\bm{\theta}}\in\Theta$, together with a loss function $\ell(f_{\bm{\theta}}(\mathbf{x}),y)$ that measures the discrepancy between the model’s output and the true label. For example, in a binary classification problem, the cross-entropy loss is given by $\ell(f_{\bm{\theta}}(\mathbf{x}),y)=-y\log\{f_{\bm{\theta}}(\mathbf{x})\}-(1-y)\log\{1-f_{\bm{\theta}}(\mathbf{x})\}$ with $f_{\bm{\theta}}(\mathbf{x})\in(0,1)$. Define $\mathcal{R}({\bm{\theta}})=\mathbb{E}\{\ell(f_{\bm{\theta}}(\mathbf{X}),Y)\}$ as the expected prediction loss, and
${\bm{\theta}}^{*}=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\mathcal{R}({\bm{\theta}}),$
as the optimal prediction parameter. The goal is to construct an estimator $\widehat{\bm{\theta}}$ such that the excess risk, $\mathcal{R}(\widehat{\bm{\theta}})-\mathcal{R}({\bm{\theta}}^{*})$, is as small as possible.
We consider a standard semi-supervised learning setting in which only a subset of the sample is labeled with the ground-truth outcome $Y$. Denote the data as $\mathcal{L}\cup\mathcal{U}$, where $\mathcal{L}=\{(\mathbf{x}_{i},y_{i}),i=1,\dots,n\}$ is labeled set and $\mathcal{U}=\{\mathbf{x}_{i},i=n+1,\dots,N\}$ is the unlabeled set. Assume that $\mathcal{L}$ and $\mathcal{U}$ are independent and share the same marginal distribution of $\mathbf{X}$.
In addition, we consider the availability of predicted labels from $K$ black-box models, denoted by $\widehat{\mathbf{Y}}=(\widehat{Y}_{1},\ldots,\widehat{Y}_{K})^{\mathrm{\scriptscriptstyle T}}$. These predictions may originate from any low-cost source: crowdsourced annotators, weak-supervision pipelines, measurement-error surrogates, model ensembles, traditional machine learning models, or large pre-trained models whose quality may be uncertain. We do not require the prediction $\widehat{Y}_{k}$ to share the same magnitude or functional form, either with one another or with the true outcome $Y$. For example, when $Y$ is a binary label, each $\widehat{Y}_{k}$ may be either categorical or continuous. We allow the generation process of $\widehat{Y}_{k}$ to be a black box, potentially depending not only on the observed feature $\mathbf{X}$ but also on some latent or unobserved variables. Predicted labels are available for all subjects, both labeled and unlabeled, and are denoted by $\widehat{\mathbf{y}}_{i}=(\widehat{y}_{1,i},\ldots,\widehat{y}_{K,i})^{\mathrm{\scriptscriptstyle T}}$, for each subject $i=1,\ldots,N$. We assume that the distribution of $\widehat{\mathbf{Y}}$ is the same for the labeled and unlabeled samples. The data structure and overall generating process are illustrated in Table 1 and Figure 1.
Table 1: Data structure in semi-supervised learning with multiple sets of predicted labels.
Unit Feature Label Multiple predicted labels
Labeled data $\mathcal{L}$ 1 $\mathbf{x}_{1}$ $y_{1}$ $\widehat{\mathbf{y}}_{1}=(\widehat{y}_{1,1},\dots,\widehat{y}_{K,1})^{\mathrm{\scriptscriptstyle T}}$
$\vdots$ $\vdots$ $\vdots$ $\quad\vdots$
$n$ $\mathbf{x}_{n}$ $y_{n}$ $\widehat{\mathbf{y}}_{n}=(\widehat{y}_{1,n},\dots,\widehat{y}_{K,n})^{\mathrm{\scriptscriptstyle T}}$
Unlabeled data $\mathcal{U}$ $n+1$ $\mathbf{x}_{n+1}$ ? $\widehat{\mathbf{y}}_{n+1}=(\widehat{y}_{1,n+1},\dots,\widehat{y}_{K,n+1})^{\mathrm{\scriptscriptstyle T}}$
$\vdots$ $\vdots$ $\vdots$ $\quad\vdots$
NN $\mathbf{x}_{N}$ ? $\widehat{\mathbf{y}}_{N}=(\widehat{y}_{1,N},\dots,\widehat{y}_{K,N})^{\mathrm{\scriptscriptstyle T}}$
3 SADA for inference
In this section, we consider the estimation and statistical inference of $\bm{\theta}$$^{*}$ defined in (1). Since the outcome is observed only in the labeled data set $\mathcal{L}$, a naive estimator, $\widehat{\bm{\theta}}^{\text{nv}}$, based solely on the labeled data, solves
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}})=\mathbf{0}.$
Assume that the Hessian matrix $\mathbf{H}=\mathbb{E}\{\partial\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}/\partial{\bm{\theta}}\}$ exists and is nonsingular. Under some regularity conditions (see Assumption 1), $\widehat{\bm{\theta}}^{\text{nv}}$ is consistent and asymptotically normal, and its mean squared error, up to a negligible term, is
$\mathbb{E}\{(\widehat{\bm{\theta}}^{\text{nv}}-{\bm{\theta}}^{*})^{\otimes 2}\}=\frac{1}{n}\mathbf{H}^{-1}{\bm{\Sigma}}_{\text{nv}}\mathbf{H}^{-1},$ (2)
where ${\bm{\Sigma}}_{\text{nv}}=\operatorname{var}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})\}$. Based on (2), valid confidence intervals for $\bm{\theta}$$^{*}$ can be constructed using $\widehat{\bm{\theta}}^{\text{nv}}$. However, since the naive estimator uses only the labeled subset of the sample, it can be substantially less efficient, especially when $n$ is small relative to $N$.
To exploit the unlabeled data, we construct a family of unbiased estimators, $\widehat{\bm{\theta}}(\mathcal{W})$, indexed by the tuning parameter $\mathcal{W}=(\mathcal{W}_{1}^{\mathrm{\scriptscriptstyle T}},\mathcal{W}_{2}^{\mathrm{\scriptscriptstyle T}},\dots,\mathcal{W}_{K}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{(Kp)\times p}$, which solves
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}})+\sum_{k=1}^{K}\mathcal{W}_{k}^{\mathrm{\scriptscriptstyle T}}\left\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathbf{s}(\mathbf{x}_{i},\widehat{y}_{k,i};{\bm{\theta}})-\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{x}_{i},\widehat{y}_{k,i};{\bm{\theta}})\right\}=\mathbf{0}.$ (3)
Noting that the augmented term has zero expectation regardless of the choice of $\mathcal{W}$ or the performance of $\widehat{\mathbf{y}}$, therefore, equation (3) is always a feasible estimating equation for $\bm{\theta}$$^{*}$. In particular, the naive estimator $\widehat{\bm{\theta}}^{\text{nv}}$ is a special case with $\mathcal{W}=\mathbf{0}$. When there is only one prediction $(K=1)$: (i) if $\mathcal{W}=\mathbf{I}$, $\widehat{\bm{\theta}}(\mathcal{W})$ reduces to the PPI estimator (Angelopoulos et al., 2023); (ii) if $\mathcal{W}=\omega\mathbf{I}$ with the tuning parameter $\omega$ selected optimally, it reduces to the PPI++ estimator (Angelopoulos et al., 2024). However, PPI and PPI++ do not provide a principled way to handle settings with multiple predictions, and even in the special case $K=1$, they do not in general achieve optimality within the family (3) in terms of the smallest possible asymptotic variance. For inference purposes, we therefore propose to identify the best estimator within the family (3) by minimizing the mean squared error of $\widehat{\bm{\theta}}(\mathcal{W})$, as formally stated in the following proposition.
Proposition 1.
Among the family of estimators $\widehat{\bm{\theta}}(\mathcal{W})$ solving (3), the optimal tuning parameter, $\mathcal{W}^{\text{opt}}$, that minimizes the mean squared error loss such that $\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W}^{\text{opt}})-{\bm{\theta}}^{*}\}^{\otimes 2}]\preceq\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W})-{\bm{\theta}}^{*}\}^{\otimes 2}]$ for any $\mathcal{W}\in\mathbb{R}^{(Kp)\times p}$ is
$\mathcal{W}^{\text{opt}}=\frac{N-n}{N}\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}^{-1}\mathbb{E}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\},$
where $\mathcal{S}(\mathbf{x},\widehat{\mathbf{y}};\theta):=(\mathbf{s}(\mathbf{x},\widehat{y}_{1};\theta)^{\mathrm{\scriptscriptstyle T}},\cdots,\mathbf{s}(\mathbf{x},\widehat{y}_{K};\theta)^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$.
The proof of Proposition 1 is given in Supplementary Section S3. Proposition 1 implies that the optimal weight $\mathcal{W}^{\text{opt}}$ exists uniquely and $\widehat{\bm{\theta}}(\mathcal{W}^{\text{opt}})$ consistently attains asymptotic variance no larger than that of the labeled-only estimator, regardless of the quality of predictions. To clarify this, one can compute that $\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W}^{\text{opt}})-{\bm{\theta}}^{*}\}^{\otimes 2}]=n^{-1}\mathbf{H}^{-1}{\bm{\Sigma}}_{\text{opt}}\mathbf{H}^{-1}$, where
${\bm{\Sigma}}_{\text{opt}}=$ $~\operatorname{var}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})\}-\frac{N-n}{N}\mathbb{E}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}$
$\qquad\qquad\qquad\qquad\qquad\times\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}^{-1}\mathbb{E}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}.$
The matrix ${\bm{\Sigma}}_{\text{opt}}$ consists of two components: the first corresponds to the variance of the naive estimator, ${\bm{\Sigma}}_{\text{nv}}$, defined in (2), and the second is a positive semi-definite matrix that represents the efficiency gain. We refer to this as a safety property, since for inference purposes, confidence intervals constructed from $\widehat{\bm{\theta}}(\mathcal{W}^{\text{opt}})$ are always more informative than those based on the naive estimator $\widehat{\bm{\theta}}^{\text{nv}}$. Beyond this safeguard, the optimal weight has another appealing adaptive feature: it assigns higher weights to more accurate predictions and downweights less accurate ones, thereby achieving greater efficiency. In particular, when some prediction is “perfect” in some sense, the procedure automatically exploits it to attain a faster convergence rate or the semiparametric efficiency. We build intuition for how the optimal weight operates and illustrate these properties using mean estimation in Section 3.1, and we formalize them in Section 3.2.
In practice, the optimal weight, $\mathcal{W}^{\text{opt}}$, can be replaced by a consistent estimator, for example,
$\widehat{\mathcal{W}}^{\text{opt}}=\bigg\{\frac{1}{N}\sum_{i=1}^{N}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};\widehat{\bm{\theta}})\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};\widehat{\bm{\theta}})^{\mathrm{\scriptscriptstyle T}}\bigg\}^{-1}\bigg\{\frac{1}{n}\sum_{i=1}^{n}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};\widehat{\bm{\theta}})\mathbf{s}(\mathbf{x}_{i},y_{i};\widehat{\bm{\theta}})^{\mathrm{\scriptscriptstyle T}}\bigg\},$
where $\widehat{\bm{\theta}}$ is a consistent estimator of $\bm{\theta}$$^{*}$ such as $\widehat{\bm{\theta}}^{\text{nv}}$. We denote the resulting estimator $\widehat{\bm{\theta}}^{\text{sada}}=\widehat{\bm{\theta}}(\widehat{\mathcal{W}}^{\text{opt}})$ as the SADA estimator. The whole protocol for calculating $\widehat{\bm{\theta}}^{\text{sada}}$ is summarized in Figure 1.
Figure 1: Protocol for computing the proposed SADA estimator $\widehat{\bm{\theta}}^{\text{sada}}$.
3.1 Illustration with mean estimation
We elaborate the intuition behind our approach using the example of mean estimation. Consider $\mathbf{s}(\mathbf{x},y;\theta)=y-\theta$, which corresponds to the estimand $\theta^{*}=\mathbb{E}(Y)$, the outcome mean. The naive estimator is the averaged outcome of the labeled samples, $\widehat{\theta}^{\text{nv}}=n^{-1}\sum_{i=1}^{n}y_{i}$. To improve efficiency by leveraging $K$ machine learning-predicted outcomes $\widehat{\mathbf{y}}_{i}$ from both labeled and unlabeled data, we consider the family of unbiased estimators $\widehat{\theta}(\bm{\omega})$, indexed by weights $\bm{\omega}=(\omega_{1},\dots,\omega_{K})^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{K}$, obtained by solving (3), namely,
$\widehat{\theta}(\bm{\omega}):=\frac{1}{n}\sum_{i=1}^{n}y_{i}+\sum_{k=1}^{K}\omega_{k}\left(\frac{1}{N-n}\sum_{i=n+1}^{N}\widehat{y}_{k,i}-\frac{1}{n}\sum_{i=1}^{n}\widehat{y}_{k,i}\right).$
As noted previously, when $\bm{\omega}=\mathbf{0}$, it reduced to the naive estimator $\widehat{\theta}^{\text{nv}}$. When there is only one prediction, i.e., $K=1$, and $\omega=1$, it turns out to be the PPI estimator (Angelopoulos et al., 2023):
$\widehat{\theta}^{\text{ppi}}=\frac{1}{n}\sum_{i=1}^{n}y_{i}+\frac{1}{N-n}\sum_{i=n+1}^{N}\widehat{y}_{i}-\frac{1}{n}\sum_{i=1}^{n}\widehat{y}_{i}.$
[…]
$\mathbb{E}[\{\widehat{\theta}(\bm{\omega})-\theta^{*}\}^{2}]=\frac{1}{n}\operatorname{var}(Y)+\underbrace{\frac{N}{n(N-n)}\bm{\omega}^{\mathrm{\scriptscriptstyle T}}\operatorname{var}(\widehat{\mathbf{Y}})\bm{\omega}-\frac{2}{n}\bm{\omega}^{\mathrm{\scriptscriptstyle T}}\operatorname{cov}(\widehat{\mathbf{Y}},Y)}_{\text{additional variance term}}.$ (4)
The first term of (4) is the variance of the naive estimator, while the second captures the variance contributed by leveraging ML predictions. Notice that (4) is a quadratic form of $\bm{\omega}$, which achieves the minimum at
$\bm{\omega}^{\text{opt}}=\frac{N-n}{N}\operatorname{var}(\widehat{\mathbf{Y}})^{-1}\operatorname{cov}(\widehat{\mathbf{Y}},Y).$ (5)
The optimal weight $\bm{\omega}^{\text{opt}}$ is fully determined by and can be easily estimated from the available data. In practical applications, one can estimate $\bm{\omega}^{\text{opt}}$ as
$\widehat{\bm{\omega}}^{\text{opt}}=\frac{N-n}{N}\bigg\{\frac{1}{N}\sum_{i=1}^{N}(\widehat{\mathbf{y}}_{i}-\overline{\widehat{\mathbf{y}}})(\widehat{\mathbf{y}}_{i}-\overline{\widehat{\mathbf{y}}})^{\mathrm{\scriptscriptstyle T}}\bigg\}^{-1}\bigg\{\frac{1}{n}\sum_{i=1}^{n}(\widehat{\mathbf{y}}_{i}-\overline{\widehat{\mathbf{y}}})(y_{i}-\overline{y})\bigg\},$
where $\overline{y}=n^{-1}\sum_{i=1}^{n}y_{i}$, and $\overline{\widehat{\mathbf{y}}}=N^{-1}\sum_{i=1}^{N}\widehat{\mathbf{y}}_{i}$. Then, the SADA estimator is given by $\widehat{\theta}^{\text{sada}}=\widehat{\theta}(\widehat{\bm{\omega}}^{\text{opt}})$. In what follows, for certain theoretical analyses and illustrations, we do not distinguish between the estimator $\widehat{\theta}(\bm{\omega}^{\text{opt}})$ and the SADA estimator $\widehat{\theta}^{\text{sada}}$, as they are asymptotically equivalent. We adopt this simplification when it does not cause confusion. We next demonstrate two key properties of the SADA estimator: safety and adaptivity.
Safety.
Based on (4) and (5), one can compute that
$\mathbb{E}\{(\widehat{\theta}^{\text{sada}}-\theta^{*})^{2}\}=\underbrace{\frac{1}{n}\operatorname{var}(Y)}_{\text{variance of }\widehat{\theta}^{\text{nv}}}-\underbrace{\left(\frac{1}{n}-\frac{1}{N}\right)\operatorname{cov}(\widehat{\mathbf{Y}},Y)^{\mathrm{\scriptscriptstyle T}}\operatorname{var}(\widehat{\mathbf{Y}})^{-1}\operatorname{cov}(\widehat{\mathbf{Y}},Y)}_{\text{efficiency gain}}.$ (6)
The efficiency gain in (6) is always non-negative regardless of the quality of the predictions. It vanishes to zero if and only if $\operatorname{cov}(\widehat{\mathbf{Y}},Y)=\mathbf{0}$; that means, none of the predictions are correlated with the ground truth $Y$. In this worst-case scenario, the optimal weight is automatically assigned to zero (i.e., $\bm{\omega}=\mathbf{0}$), reducing to the naive estimator. Except for this worst-case scenario, the SADA estimator offers a guaranteed positive efficiency gain over the naive method and enables more informative inference for $\theta^{*}$.
Adaptivity.
If any one of the predictions (no need to know which one) is highly accurate, the weight automatically picks it up and introduces an estimator with either a faster convergence rate or an improved estimation efficiency. We elaborate this in two scenarios.
First, consider the case where one of the predictions perfectly matches the ground truth, for example, $\widehat{Y}_{1}\equiv Y$. As proved in Supplementary Section S4, the optimal weight in this case is $\bm{\omega}^{\text{opt}}=(1,0,\dots,0)^{\mathrm{\scriptscriptstyle T}}\cdot(N-n)/N$. This means that the algorithm selects the most accurate prediction, $\widehat{Y}_{1}$, for estimation while discarding the less informative ones. Then, the SADA estimator turns out to be $N^{-1}\sum_{i=1}^{N}\widehat{y}_{1,i}$, and $\mathbb{E}\{(\widehat{\theta}^{\text{sada}}-\theta^{*})^{2}\}=N^{-1}\operatorname{var}(Y)$, same as the oracle estimator who knows the ground truth of the unlabeled data. In this case, the SADA estimator converges at a faster rate of $N^{-1/2}$.
Second, consider the restricted case where the predictions are deterministic functions of the available feature $\mathbf{X}$, i.e., $Y_{k}=\widehat{f}_{k}(\mathbf{X})$. Then the best prediction one can expect to fit the outcome is the conditional mean, $\mathbb{E}(Y\mid\mathbf{X})$, which minimizes the mean squared error between $Y$ and $f(\mathbf{X})$. Assume that $\widehat{Y}_{1}\equiv\mathbb{E}(Y\mid\mathbf{X})$. As proved in Supplementary Section S4, the optimal weight in this case also equals $\bm{\omega}^{\text{opt}}=(1,0,\dots,0)^{\mathrm{\scriptscriptstyle T}}\cdot(N-n)/N$. We show in the following proposition that $\widehat{\theta}^{\text{sada}}$ achieves the semiparametric efficiency bound (see Supplementary Section S1 for the definition).
Proposition 2.
Denote $r_{i}=1$ for labeled units $i=1,\dots,n$ and 0 for unlabeled units. Suppose $n/N\to\pi\in(0,1)$ as $n,N\to\infty$. Assume $\widehat{Y}_{k}=\widehat{f}_{k}(\mathbf{X})$ for $k=1,\dots,K$. Then, the efficient influence function (EIF) for estimating $\theta^{*}$, based on labeled data $\mathcal{L}=\{(\mathbf{x}_{i},y_{i}),i=1,\dots,n\}$ and unlabeled data $\mathcal{U}=\{\mathbf{x}_{i},i=n+1,\dots,N\}$, is $\phi_{\text{eif}}(r,\mathbf{x},y)=r\pi^{-1}\{y-\mathbb{E}(Y\mid\mathbf{x})\}+\mathbb{E}(Y\mid\mathbf{x})-\theta^{*}$. Suppose $\widehat{Y}_{k}\equiv\mathbb{E}(Y\mid\mathbf{X})$ for some $k$, then we have $\sqrt{N}(\widehat{\theta}^{\text{sada}}-\theta^{*})\xrightarrow{d}\mathcal{N}\{0,\mathbb{E}(\phi_{\text{eif}}^{2})\}$, which attains the semiparametric efficiency bound.
Remark 1 (Interpretation).
[…]
Remark 2 (Comparison with PPI++).
In case of the mean estimation (or, more generally, a scalar parameter $\theta^{*}$) with $K=1$, the SADA estimator $\widehat{\theta}^{\text{sada}}$ is conceptually equivalent to the PPI++ estimator (Angelopoulos et al., 2024). However, for a vector-valued parameter, the PPI++ estimator does not in general guarantee a smaller variance than the labeled-only estimator and is thus less efficient than SADA. This phenomenon is evident in our empirical results in Section 6. In addition, the PPI++ approach cannot simultaneously leverage multiple predictions when $K>1$.
3.2 Theoretical guarantees
In this section, we formally establish the theoretical properties of the SADA estimator $\widehat{\bm{\theta}}^{\text{sada}}$. We impose the following regularity conditions that ensure identification, smoothness, and well-behaved large-sample properties of the estimator.
Assumption 1.
(i) The parameter $\bm{\theta}$$^{*}$ lies in the interior of $\Theta$, and $\Theta$ is compact in $\mathbb{R}^{p}$; (ii) $\mathbb{E}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}})\}\neq 0$ if ${\bm{\theta}}\neq{\bm{\theta}}^{*}$; (iii) $\mathbf{s}(\mathbf{x},y;{\bm{\theta}})$ is differentiable with respect to $\bm{\theta}$} in a neighborhood of $\bm{\theta}$$^{*}$, and $\mathbf{H}=\mathbb{E}\{\partial\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}/\partial{\bm{\theta}}\}$ exists and nonsingular; (iv) $\mathbf{s}(\mathbf{X},Y;{\bm{\theta}})$ and $\mathbf{s}(\mathbf{X},\widehat{Y}_{k};{\bm{\theta}})$ have bounded first and second order moments in a neighborhood of $\bm{\theta}$$^{*}$ for $k=1,\dots,K$; (v) $\widehat{\mathcal{W}}^{\text{opt}}\xrightarrow{p}\mathcal{W}^{\text{opt}}$.
The following theorem formalizes the safety property of the SADA estimator and is proved in Supplementary Section S5.
Theorem 1 (Safety).
Under Assumption 1, the SADA estimator $\widehat{\bm{\theta}}^{\text{sada}}$ has the asymptotic representation $\sqrt{n}(\widehat{\bm{\theta}}^{\text{sada}}-{\bm{\theta}}^{*})\xrightarrow{d}\mathcal{N}\left(\mathbf{0},\mathbf{H}^{-1}{\bm{\Sigma}}_{\text{opt}}\mathbf{H}^{-1}\right)$. More specifically, for its mean squared error, up to a negligible term, we have
$\mathbb{E}\{(\widehat{\bm{\theta}}^{\text{sada}}-{\bm{\theta}}^{*})^{2}\}=\frac{1}{n}\mathbf{H}^{-1}{\bm{\Sigma}}_{\text{opt}}\mathbf{H}^{-1},$
where ${\bm{\Sigma}}_{\text{opt}}={\bm{\Sigma}}_{\text{nv}}-(N-n)/N\cdot{\bm{\Sigma}}_{g}$, and
${\bm{\Sigma}}_{g}=\mathbb{E}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}^{-1}\mathbb{E}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}$
is always positive semi-definite. It implies that $\operatorname{var}(\widehat{\bm{\theta}}^{\text{sada}})\preceq\operatorname{var}(\widehat{\bm{\theta}}^{\text{nv}})$.
Theorem 1 shows that the SADA estimator enjoys a guaranteed efficiency gain over the naive estimator, regardless of the quality of $\widehat{\mathbf{Y}}$. This ensures valid and more informative inference for $\bm{\theta}$$^{*}$; see Supplementary Section S2 for details. The following theorem formalizes the adaptivity property and is proved in Supplementary Section S6.
Theorem 2 (Adaptivity).
(i) Suppose $\widehat{Y}_{k}\equiv Y$ for some $k$ and Assumption 1 holds, then we have ${\bm{\Sigma}}_{g}={\bm{\Sigma}}_{\text{nv}}$, and the SADA estimator $\widehat{\bm{\theta}}^{\text{sada}}$ has the asymptotic representation $\sqrt{N}(\widehat{\bm{\theta}}^{\text{sada}}-{\bm{\theta}}^{*})\xrightarrow{d}\mathcal{N}\left(\mathbf{0},\mathbf{H}^{-1}{\bm{\Sigma}}_{\text{nv}}\mathbf{H}^{-1}\right)$. More specifically, for its mean squared error, up to a negligible term, we have
$\mathbb{E}\{(\widehat{\bm{\theta}}^{\text{sada}}-{\bm{\theta}}^{*})^{2}\}=\frac{1}{N}\mathbf{H}^{-1}{\bm{\Sigma}}_{\text{nv}}\mathbf{H}^{-1}.$
Note that this is the same as the oracle estimator who knows the ground truth of the unlabeled data; (ii) Suppose $n/N\to\pi\in(0,1)$ as $n,N\to\infty$. Assume $\widehat{Y}_{k}=\widehat{f}_{k}(\mathbf{X})$ for $k=1,\dots,K$. Then the EIF for estimating $\bm{\theta}$$^{*}$, based on labeled data $\mathcal{L}=\{(\mathbf{x}_{i},y_{i}),i=1,\dots,n\}$ and unlabeled data $\mathcal{U}=\{\mathbf{x}_{i},i=n+1,\dots,N\}$, is
$\bm{\Phi}_{\text{eif}}(r,\mathbf{x},y)=-\mathbf{H}^{-1}\left[r\pi^{-1}\{\mathbf{s}(\mathbf{x},y;{\bm{\theta}}^{*})-\bm{\mu}(\mathbf{x})\}+\bm{\mu}(\mathbf{x})\right],$
where $\bm{\mu}(\mathbf{x})=\mathbb{E}\{\mathbf{s}(\mathbf{x},Y;{\bm{\theta}}^{*})\mid\mathbf{x}\}$. Suppose $\mathbf{s}(\mathbf{x},\widehat{y}_{k^{\prime}};{\bm{\theta}}^{*})=\mathbf{s}(\mathbf{x},\widehat{f}_{k^{\prime}}(\mathbf{x});{\bm{\theta}}^{*})\equiv\bm{\mu}(\mathbf{x})$ for some $k^{\prime}$, then we have $\sqrt{N}(\widehat{\bm{\theta}}^{\text{sada}}-{\bm{\theta}}^{*})\xrightarrow{d}\mathcal{N}\{\mathbf{0},\mathbb{E}(\bm{\Phi}_{\text{eif}}\bm{\Phi}_{\text{eif}}^{\mathrm{\scriptscriptstyle T}})\}$, which attains the semiparametric efficiency bound.
Theorem 2(i) shows that, in the oracle scenario where one of the predictions coincides with the ground truth outcome, the SADA estimator effectively increases the sample size from $n$ to $N$ and thereby attains a faster convergence rate. Theorem 2(ii) shows that, when all predictions are functions of the covariates $\mathbf{X}$ and one score function equals the corresponding conditional mean, the estimator attains the semiparametric efficiency bound. Importantly, we never need to know in advance which prediction is the best; the entire procedure is data driven, which makes it appealing in practice.
4 SADA for prediction
The previous section focuses on estimation and inference for the parameter $\bm{\theta}$$^{*}$, where the primary goal is to obtain an unbiased and more efficient estimator. In other applications, such as image classification, practitioners are primarily concerned with predictive accuracy rather than parameter estimation itself, and in such cases the SADA framework needs to be adapted accordingly. Intuitively, even if one estimator, say, $\widehat{\bm{\theta}}_{1}$, is closer to $\bm{\theta}$$^{*}$ than the other estimator $\widehat{\bm{\theta}}_{2}$ in mean squared error, it does not in general imply a smaller predictive risk; that is, $\mathcal{R}(\widehat{\bm{\theta}}_{1})\leq\mathcal{R}(\widehat{\bm{\theta}}_{2})$. Motivated by this distinction, in this section we slightly modify the SADA procedure to explicitly target minimization of the predictive risk $\mathcal{R}({\bm{\theta}})$ and discuss how it can be designed to achieve better prediction performance.
To build intuition, consider an unbiased estimator $\widehat{\mathcal{R}}({\bm{\theta}})$ of $\mathcal{R}({\bm{\theta}})$ such that $\mathbb{E}\{\widehat{\mathcal{R}}({\bm{\theta}})\}=\mathcal{R}({\bm{\theta}})$. Define $\widehat{\bm{\theta}}=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\widehat{\mathcal{R}}({\bm{\theta}})$. Then, in a heuristic sense, the excess risk $\mathcal{R}(\widehat{\bm{\theta}})-\mathcal{R}({\bm{\theta}}^{*})\leq\mathcal{R}(\widehat{\bm{\theta}})-\widehat{\mathcal{R}}(\widehat{\bm{\theta}})+\widehat{\mathcal{R}}({\bm{\theta}}^{*})-\mathcal{R}({\bm{\theta}}^{*})\leq 2\sup_{{\bm{\theta}}\in\Theta}|\widehat{\mathcal{R}}({\bm{\theta}})-\mathcal{R}({\bm{\theta}})|$, where the first inequality follows from $\widehat{\mathcal{R}}(\widehat{\bm{\theta}})\leq\widehat{\mathcal{R}}({\bm{\theta}}^{*})$ by definition. Under some regularity conditions, standard uniform concentration results indicate that this supremum depends on the complexity of $\Theta$, the variance scale of $\widehat{\mathcal{R}}$, and the tail behavior of the empirical loss. Accordingly, once class complexity is fixed, reducing the variance of $\widehat{\mathcal{R}}$ is a direct and effective way to tighten the uniform deviation and lower the excess risk.
Let $\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})=n^{-1}\sum_{i=1}^{n}\ell(f_{\bm{\theta}}(\mathbf{x}_{i}),y_{i})$ denote the naive empirical loss function which only uses the labeled data, and $\widehat{\bm{\theta}}^{\text{nv}}=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})$. We consider a set of empirical loss functions, indexed by $\bm{\omega}=(w_{1},w_{2},\dots,w_{K})^{\mathrm{\scriptscriptstyle T}}$,
$\widehat{\mathcal{R}}({\bm{\theta}},\bm{\omega})=\frac{1}{n}\sum_{i=1}^{n}\ell(f_{\bm{\theta}}(\mathbf{x}_{i}),y_{i})+\bm{\omega}^{\mathrm{\scriptscriptstyle T}}\left\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathcal{L}(f_{\bm{\theta}}(\mathbf{x}_{i}),\widehat{\mathbf{y}}_{i})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(f_{\bm{\theta}}(\mathbf{x}_{i}),\widehat{\mathbf{y}}_{i})\right\},$ (7)
where $\mathcal{L}(f_{\bm{\theta}}(\mathbf{x}),\widehat{\mathbf{y}})=(\ell(f_{\bm{\theta}}(\mathbf{x}),\widehat{y}_{1}),\ell(f_{\bm{\theta}}(\mathbf{x}),\widehat{y}_{2}),\dots,\ell(f_{\bm{\theta}}(\mathbf{x}),\widehat{y}_{K}))^{\mathrm{\scriptscriptstyle T}}$. Noting that the second term has zero expectation for any $\bm{\omega}$ and any distribution of $\widehat{\mathbf{y}}$, therefore, equation (7) always defines an unbiased estimator for $\mathcal{R}({\bm{\theta}})$. As a special case, when $\bm{\omega}=\mathbf{0}$, $\widehat{\mathcal{R}}({\bm{\theta}},\bm{\omega})$ reduces to $\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})$. For a fixed $\bm{\theta}$}, we propose to select the optimal loss function from the set by minimizing the mean squared error between $\widehat{\mathcal{R}}({\bm{\theta}},\bm{\omega})$ and the expected loss $\mathcal{R}({\bm{\theta}})$. We have
$\bm{\omega}^{\text{opt}}({\bm{\theta}})=$ $~\operatornamewithlimits{arg\,min}_{\bm{\omega}}\mathbb{E}\{\widehat{\mathcal{R}}({\bm{\theta}},\bm{\omega})-\mathcal{R}({\bm{\theta}})\}^{2}$
$=$ $~\frac{N-n}{N}\operatorname{var}\{\mathcal{L}(f_{\bm{\theta}}(\mathbf{X}),\widehat{\mathbf{Y}})\}^{-1}\operatorname{cov}\{\mathcal{L}(f_{\bm{\theta}}(\mathbf{X}),\widehat{\mathbf{Y}}),\ell(f_{\bm{\theta}}(\mathbf{X}),Y)\}.$
Then, we define $\widehat{\mathcal{R}}^{\text{sada}}({\bm{\theta}})=\widehat{\mathcal{R}}({\bm{\theta}},\widehat{\bm{\omega}}^{\text{opt}}({\bm{\theta}}))$, where $\widehat{\bm{\omega}}^{\text{opt}}$ is a consistent estimator of $\bm{\omega}^{\text{opt}}$, and obtain the SADA estimator as $\widehat{\bm{\theta}}^{\text{sada}}=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\widehat{\mathcal{R}}^{\text{sada}}({\bm{\theta}})$. It is worth noting that the optimal weight depends on $\bm{\theta}$} and, in principle, should be optimized jointly with the parameter within the loss function, which may lead to numerical difficulties and instability in practice. Therefore, for implementation, we recommend an iterative procedure that alternately updates the weight $\bm{\omega}^{\text{opt}}$ and the target parameter $\bm{\theta}$$^{*}$, as outlined in Algorithm 1.
Algorithm 1 SADA for prediction
1:Input: Observations $(\mathbf{x}_{i},y_{i},\widehat{\mathbf{y}}_{i})_{i=1}^{n}\cup(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i})_{i=n+1}^{N}$, prediction function $f_{\bm{\theta}}$, loss function $\ell(f_{\bm{\theta}}(\mathbf{x}),y)$, number of iterations $J$.
2: Compute the naive estimator $\widehat{\bm{\theta}}^{\text{nv}}=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}}n^{-1}\sum_{i=1}^{n}\ell(f_{\bm{\theta}}(\mathbf{x}_{i}),y_{i})$.
3: Initialize $\widehat{\bm{\theta}}^{(0)}=\widehat{\bm{\theta}}^{\text{nv}}$.
4:for $j=1,\ldots,J$ do
5:  Compute the optimal weight
$\widehat{\bm{\omega}}^{(j)}=\frac{N-n}{N}$ $\Bigg\{\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}\left(f_{\widehat{\bm{\theta}}^{(j-1)}}(\mathbf{x}_{i}),\widehat{\mathbf{y}}_{i}\right)\mathcal{L}\left(f_{\widehat{\bm{\theta}}^{(j-1)}}(\mathbf{x}_{i}),\widehat{\mathbf{y}}_{i}\right)^{\mathrm{\scriptscriptstyle T}}\Bigg\}^{-1}$
$~~\times\Bigg\{\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}\left(f_{\widehat{\bm{\theta}}^{(j-1)}}(\mathbf{x}_{i}),\widehat{\mathbf{y}}_{i}\right)\ell\left(f_{\widehat{\bm{\theta}}^{(j-1)}}(\mathbf{x}_{i}),y_{i}\right)\Bigg\}.$
6:  Update the parameter estimate
$\widehat{\bm{\theta}}^{(j)}=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\ell(f_{\bm{\theta}}(\mathbf{x}_{i}),y_{i})+\Bigg\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathcal{L}(f_{\bm{\theta}}(\mathbf{x}_{i}),\widehat{\mathbf{y}}_{i})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(f_{\bm{\theta}}(\mathbf{x}_{i}),\widehat{\mathbf{y}}_{i})\Bigg\}^{\mathrm{\scriptscriptstyle T}}\widehat{\bm{\omega}}^{(j)}.$
7:end for
8:Output: SADA estimator $\widehat{\bm{\theta}}^{\text{sada}}=\widehat{\bm{\theta}}^{(J)}$.
4.1 Theoretical properties
In this section, we investigate the theoretical properties of $\widehat{\bm{\theta}}^{\text{sada}}$ in terms of its excess risk, $\mathcal{R}(\widehat{\bm{\theta}}^{\text{sada}})-\mathcal{R}({\bm{\theta}}^{*})$. We first introduce some notations. For a metric space $(\Theta,d)$, let $N(\epsilon,\Theta,d)$ denote the covering number, i.e., the minimum number of balls of radius $\epsilon$ with respect to the metric $d$ that cover the space $\Theta$. For a positive constant $C$, define a class of sub-Gaussian random variables
$\mathcal{F}_{C}:=\left\{\xi:\text{$\xi$ is $\sigma$-sub-Gaussian for some }\sigma\leq C\sqrt{\operatorname{var}(\xi)}\right\}.$
We use some simplified notations when no confusion arises. Let $\ell_{\bm{\theta}}=\ell(f_{\bm{\theta}}(\mathbf{X}),\mathbf{Y})$, $\widetilde{\ell}_{\bm{\theta}}=\ell_{\bm{\theta}}-\mathcal{L}(f_{\bm{\theta}}(\mathbf{X}),\widehat{\mathbf{Y}})^{\mathrm{\scriptscriptstyle T}}\bm{\omega}^{\text{opt}}({\bm{\theta}})$, and $\widehat{\ell}_{\bm{\theta}}=N/(N-n)(\ell_{\bm{\theta}}-\widetilde{\ell}_{\bm{\theta}})$. We endow the parameter space $\Theta$ with three metrics: $d_{1}({\bm{\theta}},{\bm{\theta}}^{\prime})=\operatorname{sd}(\ell_{\bm{\theta}}-\ell_{{\bm{\theta}}^{\prime}})$, $d_{2}({\bm{\theta}},{\bm{\theta}}^{\prime})=\operatorname{sd}(\widehat{\ell}_{\bm{\theta}}-\widehat{\ell}_{{\bm{\theta}}^{\prime}})$ and $d_{3}({\bm{\theta}},{\bm{\theta}}^{\prime})=\operatorname{sd}(\widetilde{\ell}_{\bm{\theta}}-\widetilde{\ell}_{{\bm{\theta}}^{\prime}})$, where $\operatorname{sd}(U)=\sqrt{\operatorname{var}(U)}$ denotes the standard deviation for a random variable $U$. We impose the following regularity conditions, which control the smoothness and tail behavior of the loss functions and the complexity of the parameter space.
Assumption 2.
(i) The loss function $\ell(f_{\bm{\theta}}(\mathbf{x}),y)$ is continuous in $\bm{\theta}$} and uniformly bounded for all ${\bm{\theta}}\in\Theta$ and $(x,y)\in\mathcal{X}\times\mathcal{Y}$; (ii) There exists a positive constant $C>0$ such that, $\forall{\bm{\theta}},{\bm{\theta}}^{\prime}\in\Theta$, $\ell_{\bm{\theta}}-\ell_{{\bm{\theta}}^{\prime}}$, $\widehat{\ell}_{\bm{\theta}}-\widehat{\ell}_{{\bm{\theta}}^{\prime}}$ and $\widetilde{\ell}_{\bm{\theta}}-\widetilde{\ell}_{{\bm{\theta}}^{\prime}}$ all belong to $\mathcal{F}_{C}$. (iii) The Dudley’s entropy integral is finite for $d_{1},d_{2}$ and $d_{3}$; that is, for $j=1,2,3,$
$\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{j})}d\epsilon<\infty.$
Theorem 3 (Excess risk).
Under Assumption 2, we have that with probability at least $1-\delta$,
$\mathcal{R}(\widehat{\bm{\theta}}^{\text{sada}})-\mathcal{R}({\bm{\theta}}^{*})\leq\left\{\begin{aligned} &\mathcal{O}\left(\sqrt{\frac{1}{N}\log\frac{1}{\delta}}+\sqrt{\frac{1}{N}}\right),~~\text{if $\widehat{Y}_{k}=Y$ for some $k\in\{1,\dots,K\}$},\\ &\mathcal{O}\left(\sqrt{\frac{1}{n}\log\frac{1}{\delta}}+\sqrt{\frac{1}{n}}+\sqrt{\frac{1}{N}\log\frac{1}{\delta}}+\sqrt{\frac{1}{N}}\right),~~\text{otherwise.}\end{aligned}\right$.
A more precise non-asymptotic bound for $\mathcal{R}(\widehat{\bm{\theta}}^{\text{sada}})-\mathcal{R}({\bm{\theta}}^{*})$ is given in Theorem S4 in Supplementary Section S7. As a comparison, Theorem S4 also shows that the excess risk of the naive estimator, $\mathcal{R}(\widehat{\bm{\theta}}^{\text{nv}})-\mathcal{R}({\bm{\theta}}^{*})$, converges at the order $\mathcal{O}(\sqrt{n^{-1}\log(\delta^{-1})}+\sqrt{n^{-1}})$ with high probability. This implies that the excess risk of the SADA estimator converges at a rate no slower than that of the labeled data only baseline, regardless of the quality of the predictions. In contrast, if one of the predictions perfectly fits the ground truth, although it is unknown which one, the SADA estimator achieves an even faster convergence rate, effectively increasing the sample size from $n$ to $N$.
5 Simulation studies
In this section, we conduct some simulation studies to evaluate the performance of the proposed method. We generate $Y\sim\mathcal{N}(\theta^{*},1)$ with $\theta^{*}=0.5$, and generate two predictions $\widehat{Y}_{1}=\gamma Y+(1-\gamma)\epsilon_{1}$ and $\widehat{Y}_{2}=(1-\gamma)Y+\gamma\epsilon_{2}$, where $\gamma$ is a tuning parameter and $\epsilon_{1},\epsilon_{2}$ are independent white noises drawn from the standard normal distribution. The goal is to estimate and to conduct inference for $\theta^{*}$. As $\gamma$ increases from 0 to 1, the quality of the prediction $\widehat{Y}_{1}$ improves in approximating the ground truth $Y$, while the prediction $\widehat{Y}_{2}$ becomes less accurate. We generate a dataset of size $N=200$, among which $n=60$ observations are labeled, and repeat the experiment over 1000 Monte Carlo replications.
We include naive, PPI, and PPI++ methods for comparison. Table 2 reports the relative efficiency of each method relative to the naive method, defined as the ratio of their standard deviations across replications as $\gamma$ varies from 0 to 1, and Figure 2 presents the corresponding plots. The results on bias and coverage probability are reported in Supplementary Section S8, as they show negligible differences across methods and are not the primary criteria for assessing their performance. It is shown that, the performance of the PPI estimator is highly sensitive to prediction accuracy: it can perform even worse than the naive estimator when the prediction is poor, as illustrated in Figure 2(a), where the ratios of standard deviations exceed 1 when $\gamma$ approaches 0 for PPI using $\widehat{Y}_{1}$ and 1 for PPI using $\widehat{Y}_{2}$. By contrast, PPI++ safeguards against low quality predictions, achieving a variance that is never larger than that of the naive estimator, as shown in Figure 2(b). However, its efficiency gains arise only when the predictions are informative, and it reduces to the naive estimator when the predictions are noninformative.
In comparison, the SADA estimator exhibits several favorable properties. First, it consistently attains a smaller standard deviation than the naive estimator across the entire range of prediction quality. Second, it adaptively exploits the relative strengths of $\widehat{Y}_{1}$ and $\widehat{Y}_{2}$ in the regions where each is more informative, which leads to the most stable performance among all methods considered. Third, it is able to incorporate information from both predictions through data-driven weights. This is particularly evident at $\gamma=0.5$, where SADA has substantially lower standard deviations than PPI++ under either prediction.
Table 2: Relative efficiency (ratio of standard deviations for each method to that of the naive method) under varying prediction quality $\gamma$.
Methods Relative efficiency
$\gamma=0$ $\gamma=0.1$ $\gamma=0.2$ $\gamma=0.3$ $\gamma=0.4$ $\gamma=0.5$ $\gamma=0.6$ $\gamma=0.7$ $\gamma=0.8$ $\gamma=0.9$ $\gamma=1$
PPI $\widehat{Y}_{1}$ 1.549 1.396 1.247 1.104 0.970 0.842 0.736 0.653 0.608 0.610 0.654
$\widehat{Y}_{2}$ 0.655 0.609 0.611 0.656 0.737 0.845 0.968 1.103 1.245 1.392 1.551
PPI++ $\widehat{Y}_{1}$ 0.989 0.982 0.965 0.933 0.877 0.799 0.713 0.636 0.582 0.556 0.549
$\widehat{Y}_{2}$ 0.549 0.556 0.585 0.639 0.715 0.801 0.876 0.932 0.965 0.982 0.989
SADA both 0.548 0.555 0.584 0.636 0.697 0.727 0.697 0.635 0.583 0.555 0.548
(a) PPI
(b) PPI++
(c) SADA (ours)
Figure 2: Relative efficiency of different methods compared to the naive method as prediction quality varies. Stars and triangles on the line indicate standard deviations over 1000 replications; scatter points represent those from individual replications.
6 Real data applications
In this section, we apply the proposed methods to two real-data analyses with distinct tasks. The first is a regression setting based on online requests from Stack Exchange and Wikipedia (Danescu-Niculescu-Mizil et al., 2013), where we conduct estimation and inference on the regression coefficient linking perceived politeness to specific linguistic features. The second is an image classification task based on the ImageNet dataset (Russakovsky et al., 2015).
6.1 Politeness of online requests
We study how certain linguistic features affect perceived politeness using the dataset from Danescu-Niculescu-Mizil et al. (2013). The data consist of over ten thousand two-sentence requests collected from Wikipedia and Stack Exchange. Each request was independently annotated by five U.S.-based Amazon Mechanical Turk workers on a 1-25 scale, and the final politeness score is the average of these five ratings. The dataset also includes automatically extracted politeness markers such as gratitude, deference, greetings, sentiment cues, modality choices, and pronominal patterns, obtained by combining dependency parsing with curated lexical resources. In this experiment, we focus on the relationship between politeness and indicative modal features. We select 1000 requests from the dataset and treat the regression coefficient from regressing the politeness score on the indicative modal features as the parameter of interest. Among these requests, we randomly designate 300 as unlabeled data and vary the number of labeled samples $n$ from 50 to 700.
To evaluate the proposed methods, we also obtain predictions for each request from three large language models, GPT-4o, Llama-3-8B, and DeepSeek-V3, by prompting each model to rate the politeness of the given text on a scale of 1-25. For each request, we supply the model with the text under review, relevant background information, and a set of ten independent prompts designed to reduce randomness in the generated scores. The final prediction is computed as the average of the ten outputs. Appendix S9 provides a step-by-step description of the procedure used to generate these predictions.
Figure 3 reports the variation of standard deviations of different methods as the labeled data size increases. The three panels differ only in the implementation of the PPI and PPI++ methods, with each panel using predictions from a different single LLM. While the standard deviations of all methods generally decrease with more labeled data, the PPI and PPI++ estimators do not consistently improve upon the naive approach in terms of efficiency. In contrast, the SADA estimator, which integrates all predictions, consistently outperforms all other methods. Its performance closely tracks that of the best-performing PPI++ estimator based on GPT-4o, highlighting its adaptive behavior. We also plot the oracle SADA estimator, SADAo, which includes the ground truth as one of its predictions. As shown, SADAo performs almost identically to the oracle benchmark obtained by regressing the ground-truth outcome on the covariates using the full data, further confirming the desired adaptivity of the proposed method.
Figure 3: Comparison of standard deviations of different methods leveraging various prediction strategies. The estimand is the regression coefficient of politeness score on indicative modal features.
6.2 Image Classification
In this section, for prediction purposes, we apply SADA to the ImageNet-5 data set, which is constructed from a random subset of five classes from ImageNet-1k (Russakovsky et al., 2015). The data consist of five visually diverse object categories, including animals, artifacts, and natural scenes, each paired with a ground truth class label from the human annotated ImageNet hierarchy. The data set contains 1300 images per class, for a total of 6500 images across five classes, and is split into 5000 training images, 750 validation images, and 750 test images. Among the training images, we randomly select 2500 as unlabeled data and vary the labeled-unlabeled ratio from 0.2 to 1.
We generate two pseudo labels $\widehat{Y}_{1}$ and $\widehat{Y}_{2}$ by training two teacher models, EfficientNet-B0 (Tan & Le, 2020) and RegNetY-008 (Radosavovic et al., 2020), using the following hyperparameters: 10 training epochs, a learning rate of 1e-3, a batch size of 256, and the AdamW optimizer. These trained models are then used to generate pseudo labels for all 5000 training images.
We evaluate the different methods by training DaViT-T (Ding et al., 2022), a popular vision transformer architecture that achieves state-of-the-art performance on ImageNet. The model is trained for 15 epochs with a learning rate of 1e-3, a batch size of 256, and the default AdamW optimizer. Figure 4 reports the Top-1 accuracy on the test set for the naive baseline, PPI using $\widehat{Y}_{1}$ and $\widehat{Y}_{2}$ separately, and SADA across labeled-unlabeled ratios ranging from 0.2 to 1. It is shown that SADA consistently achieves the best performance in terms of prediction accuracy across all settings.
Figure 4: Top-1 accuracy of various methods on the test images across increasing numbers of labeled data.
7 Discussion
In this paper, we propose a novel procedure, SADA, that safely and adaptively aggregates multiple predicted labels originating from diverse state-of-the-art machine learning models for both inference and prediction purposes. An attractive feature of SADA is that it does not require any assumptions on the form, quality, or source of the predictions, which may range from highly accurate and close to the ground truth to extremely poor and essentially random noise. The algorithm automatically aggregates these predictions by assigning them appropriate weights, leading to more efficient inference or improved predictive accuracy. In addition, SADA is practitioner friendly, as it does not require tuning of hyperparameters or any preliminary assessment of the quality of the available predictions.
There are several directions for future exploration. An essential assumption in this paper is that the labeled and unlabeled data are drawn from the same distribution. In settings with distribution shift, the current method should be adapted accordingly, because the augmentation terms in (3) will not converge to zero. In such cases, distributional adjustment based on the type of shift, such as covariate shift or label shift between the labeled and unlabeled data, is necessary to restore the effectiveness of the proposed methods. We are going to investigate these extensions in future work.
Supplementary material
The supplementary material contains all technical proofs as well as additional experimental details. The R package, sada, for implementing the proposed algorithms is available at https://github.com/jw-shan/sada.
SUPPLEMENT
Appendix S1 Semiparametric efficiency bound
Here we briefly introduce the regular and asymptotically linear (RAL) estimator and influence function. In general, given i.i.d. copies of the random sample $\left\{\mathbf{z}_{1},\ldots,\mathbf{z}_{n}\right\}$ with sample size $n$, an estimator for the parameter of interest $\bm{\beta}$, $\widehat{\bm{\beta}}$, is a RAL estimator if
$\sqrt{n}(\widehat{\bm{\beta}}-\bm{\beta})=n^{-1/2}\sum_{i=1}^{n}\bm{\phi}\left(\mathbf{z}_{i}\right)+o_{p}(1),$
where the zero-mean function $\bm{\phi}(\cdot)$ is called the influence function of $\widehat{\bm{\beta}}$. Then, the central limit theorem implies that $\sqrt{n}(\widehat{\bm{\beta}}-\bm{\beta})\xrightarrow{d}\mathcal{N}(\mathbf{0},\mathbb{E}(\bm{\phi}\bm{\phi}^{\mathrm{\scriptscriptstyle T}}))$, provided that $\mathbb{E}(\bm{\phi}\bm{\phi}^{\mathrm{\scriptscriptstyle T}})$ is finite and nonsingular. Among all RAL estimators for $\bm{\beta}$, the influence function of the one with the smallest asymptotic variance is called the efficient influence function (EIF), $\bm{\phi}_{\text{eif}}$, and the semiparametric efficiency bound is $\mathbb{E}\left(\bm{\phi}_{\text{eif}}\bm{\phi}_{\text{eif}}^{\mathrm{\scriptscriptstyle T}}\right)$.
We provide a brief, non-technical overview of the derivation of influence functions. At first, the likelihood can be decomposed into a parametric component and one or more nonparametric components. For each nonparametric component, we can identify a corresponding nuisance tangent space. The overall nuisance tangent space for the semiparametric model is then obtained by combining the individual nuisance tangent spaces. A valid influence function must lie in the orthogonal complement of this nuisance tangent space. Under suitable regularity conditions, an element from this perpendicular space can be chosen as the influence function.
We refer readers to Bickel et al. (1993) and Tsiatis (2006) for further interpretations. In addition, Kennedy (2023) provided several strategies for deriving the efficient influence function together with many examples.
Appendix S2 Procedure of inference
The asymptotic normality of $\widehat{\bm{\theta}}^{\text{sada}}$ established in Theorem 1 enables us to construct confidence intervals for $\bm{\theta}$$^{*}$ following standard procedure. Specifically, let $\widehat{\mathbf{H}}=n^{-1}\sum_{i=1}^{n}\partial\mathbf{s}(\mathbf{X}_{i},Y_{i};\widehat{\bm{\theta}}^{\text{sada}})/\partial{\bm{\theta}}$, $\widehat{\bm{\Sigma}}_{\text{opt}}=\widehat{\bm{\Sigma}}_{\text{nv}}-(N-n)/N\cdot\widehat{\bm{\Sigma}}_{g}$, where
$\widehat{\bm{\Sigma}}_{\text{nv}}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{X}_{i},Y_{i};\widehat{\bm{\theta}}^{\text{sada}})\mathbf{s}(\mathbf{X}_{i},Y_{i};\widehat{\bm{\theta}}^{\text{sada}})^{\mathrm{\scriptscriptstyle T}},$
and
$\widehat{\bm{\Sigma}}_{g}=$ $\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{X}_{i},Y_{i};\widehat{\bm{\theta}}^{\text{sada}})\mathcal{S}(\mathbf{X}_{i},\widehat{\mathbf{Y}}_{i};\widehat{\bm{\theta}}^{\text{sada}})^{\mathrm{\scriptscriptstyle T}}\right\}$
$\times\left\{\frac{1}{N}\sum_{i=1}^{N}\mathcal{S}(\mathbf{X}_{i},\widehat{\mathbf{Y}}_{i};\widehat{\bm{\theta}}^{\text{sada}})\mathcal{S}(\mathbf{X}_{i},\widehat{\mathbf{Y}}_{i};\widehat{\bm{\theta}}^{\text{sada}})^{\mathrm{\scriptscriptstyle T}}\right\}^{-1}$
$\times\left\{\frac{1}{n}\sum_{i=1}^{n}\mathcal{S}(\mathbf{X}_{i},\widehat{\mathbf{Y}}_{i};\widehat{\bm{\theta}}^{\text{sada}})\mathbf{s}(\mathbf{X}_{i},Y_{i};\widehat{\bm{\theta}}^{\text{sada}})^{\mathrm{\scriptscriptstyle T}}\right\}.$
Let $\widehat{\bm{\Omega}}=\widehat{\mathbf{H}}^{-1}\widehat{\bm{\Sigma}}_{\text{opt}}\widehat{\mathbf{H}}^{-1}$. Then, a confidence region $C_{\alpha}$ of $\bm{\theta}$$^{*}$ with error level $\alpha$ is given by
$C_{\alpha}=\{{\bm{\theta}}\text{ such that }\sqrt{n}\widehat{\bm{\Omega}}^{-1/2}(\widehat{\bm{\theta}}^{\text{sada}}-{\bm{\theta}})\in\mathcal{R}_{\alpha}\},$
where $\mathcal{R}_{\alpha}$ is a region in $\mathbb{R}^{p}$ such that $\mathbb{P}(\bm{Z}_{p}\in\mathcal{R}_{\alpha})=1-\alpha$ for a $p$-dimensional standard normal random vector $\mathbf{Z}_{p}$. For point-wise inference of $j$th component of $\bm{\theta}$$^{*}$, $\theta_{j}^{*}$, let $\widehat{\bm{\Omega}}_{jj}$ be the $j$th diagonal entry of $\widehat{\bm{\Omega}}$. Then, a confidence interval $C_{j,\alpha}$ of $\bm{\theta}$$^{*}$ with error level $\alpha$ is given by
$C_{j,\alpha}=\widehat{\theta}_{j}^{\text{sada}}\pm\widehat{\bm{\Omega}}_{jj}^{1/2}Z_{1-\alpha/2}/\sqrt{n},$
where $Z_{1-\alpha/2}$ is the $1-\alpha/2$ upper quantile of the standard normal distribution.
Appendix S3 Proof of Proposition 1
Proof.
Recall that $\widehat{\bm{\theta}}(\mathcal{W})$ solves
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}})+\mathcal{W}^{\mathrm{\scriptscriptstyle T}}\left\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}})\right\}=\mathbf{0},$
where $\mathcal{W}=(\mathcal{W}_{1}^{\mathrm{\scriptscriptstyle T}},\mathcal{W}_{2}^{\mathrm{\scriptscriptstyle T}},\dots,\mathcal{W}_{K}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{(Kp)\times p}$, and $\mathcal{S}(\mathbf{x},\widehat{\mathbf{y}};{\bm{\theta}})=(\mathbf{s}(\mathbf{x},\widehat{y}_{1};{\bm{\theta}})^{\mathrm{\scriptscriptstyle T}},\cdots,\mathbf{s}(\mathbf{x},\widehat{y}_{K};{\bm{\theta}})^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. Then the standard Taylor expansion yields that
$\mathbf{0}=$ $~\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}}^{*})+\mathcal{W}^{\mathrm{\scriptscriptstyle T}}\left\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})\right\}$
$~+\left[\frac{1}{n}\sum_{i=1}^{n}\frac{\partial}{\partial{\bm{\theta}}^{\mathrm{\scriptscriptstyle T}}}\mathbf{s}(\mathbf{X},Y;\overline{{\bm{\theta}}})+\mathcal{W}^{\mathrm{\scriptscriptstyle T}}\left\{\frac{1}{N-n}\sum_{i=n+1}^{N}\frac{\partial}{\partial{\bm{\theta}}^{\mathrm{\scriptscriptstyle T}}}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};\overline{{\bm{\theta}}})-\frac{1}{n}\sum_{i=1}^{n}\frac{\partial}{\partial{\bm{\theta}}^{\mathrm{\scriptscriptstyle T}}}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};\overline{{\bm{\theta}}})\right\}\right]$
$\hskip 327.20668pt\times\{\widehat{\bm{\theta}}(\mathcal{W})-{\bm{\theta}}^{*}\},$
where $\overline{{\bm{\theta}}}$ lies between $\widehat{\bm{\theta}}(\mathcal{W})$ and $\bm{\theta}$$^{*}$. Then following uniform weak law of large number (Newey & McFadden 1994), under the regularity conditions, we have
$\widehat{\bm{\theta}}(\mathcal{W})-{\bm{\theta}}^{*}\doteq-\mathbf{H}^{-1}\left[\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}}^{*})+\mathcal{W}^{\mathrm{\scriptscriptstyle T}}\left\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})\right\}\right],$
where $\mathbf{H}=\mathbb{E}\{\partial\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}/\partial{\bm{\theta}}\}$. Then
$\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W})$- ${\bm{\theta}}^{*}\}^{\otimes 2}]\doteq\mathbf{H}^{-1}\Bigg[\frac{1}{n}\operatorname{var}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})\}+\frac{N}{n(N-n)}\mathcal{W}^{\mathrm{\scriptscriptstyle T}}\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}\mathcal{W}$
$-\frac{1}{n}\mathbb{E}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}\mathcal{W}-\frac{1}{n}\mathcal{W}^{\mathrm{\scriptscriptstyle T}}\mathbb{E}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}\Bigg]\mathbf{H}^{-1}.$
Let
$\mathcal{W}^{\text{opt}}=\frac{N-n}{N}\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}^{-1}\mathbb{E}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}.$
We next show that
$\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W}^{\text{opt}})-{\bm{\theta}}^{*}\}^{\otimes 2}]\preceq\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W})-{\bm{\theta}}^{*}\}^{\otimes 2}],~~\forall\mathcal{W}\in\mathbb{R}^{(Kp)\times p},$ (S1)
i.e., $\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W})-{\bm{\theta}}^{*}\}^{\otimes 2}]-\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W}^{\text{opt}})-{\bm{\theta}}^{*}\}^{\otimes 2}]$ is a positive semi-definite matrix for any $\mathcal{W}\in\mathbb{R}^{(Kp)\times p}$. Note that
$\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W}^{\text{opt}})-{\bm{\theta}}^{*}\}^{\otimes 2}]=$ $~\mathbf{H}^{-1}\Bigg[\frac{1}{n}\operatorname{var}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})\}-\left(\frac{1}{n}-\frac{1}{N}\right)\mathbb{E}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}$
$\qquad\qquad\times\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}^{-1}\mathbb{E}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}\Bigg]\mathbf{H}^{-1}.$
For any non-zero vector $\mathbf{a}\in\mathbb{R}^{p}$, let $\widetilde{\mathbf{a}}=\mathbf{H}^{-1}\mathbf{a}$, then we have
$\mathbf{a}^{\mathrm{\scriptscriptstyle T}}\left(\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W})-{\bm{\theta}}^{*}\}^{\otimes 2}]-\mathbb{E}[\{\widehat{\bm{\theta}}(\mathcal{W}^{\text{opt}})-{\bm{\theta}}^{*}\}^{\otimes 2}]\right)\mathbf{a}$
$=$ $~\frac{N}{n(N-n)}\left(\mathcal{W}\widetilde{\mathbf{a}}-\frac{N-n}{N}\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}^{-1}\mathbb{E}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}\right)^{\mathrm{\scriptscriptstyle T}}$
$\qquad\qquad\times\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}$
$\qquad\qquad\times\left(\mathcal{W}\widetilde{\mathbf{a}}-\frac{N-n}{N}\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}^{-1}\mathbb{E}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}\right)$
$\geq$ $~0.$
Therefore, (S1) holds by definition. ∎
Appendix S4 Proofs of adaptivity for mean estimation
We provide the proofs for the two cases of the adaptivity properties in the illustrative mean estimation example described in Section 3.1.
Proof of the oracle case.
Assume $\widehat{Y}_{1}\equiv Y$. We denote $\widehat{\mathbf{Y}}_{-1}=(\widehat{\mathbf{Y}}_{2},\dots,\widehat{\mathbf{Y}}_{K})^{\mathrm{\scriptscriptstyle T}}$, and
$\operatorname{var}(\widehat{\mathbf{Y}})=\left(\begin{array}[]{cc}\operatorname{var}(Y)&\operatorname{cov}(\widehat{\mathbf{Y}}_{-1},Y)^{\mathrm{\scriptscriptstyle T}}\\ \operatorname{cov}(\widehat{\mathbf{Y}}_{-1},Y)&\operatorname{var}(\widehat{\mathbf{Y}}_{-1})\end{array}\right)=:\left(\begin{array}[]{cc}C_{11}&\bm{C}_{21}^{\mathrm{\scriptscriptstyle T}}\\ \bm{C}_{21}&\bm{C}_{22}\end{array}\right).$
By the inversion formula of $2\times 2$ block matrix (Lu & Shiou 2002, Theorem 2.1), we have
$\operatorname{var}(\widehat{\mathbf{Y}})^{-1}=\left(\begin{matrix}C_{11}^{-1}+C_{11}^{-1}\bm{C}_{21}\bm{M}^{-1}\bm{C}_{21}^{\mathrm{\scriptscriptstyle T}}C_{11}^{-1}&-C_{11}^{-1}\bm{C}_{21}\bm{M}^{-1}\\ -\bm{M}^{-1}\bm{C}_{21}^{\mathrm{\scriptscriptstyle T}}C_{11}^{-1}&\bm{M}^{-1}\end{matrix}\right),$ (S2)
and $\operatorname{cov}(\widehat{\mathbf{Y}},Y)=(C_{11},\bm{C}_{21}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$, where $\bm{M}=\bm{C}_{22}-\bm{C}_{21}^{\mathrm{\scriptscriptstyle T}}C_{11}^{-1}\bm{C}_{21}$. Then
$\bm{\omega}^{\text{opt}}=\frac{N-n}{N}\operatorname{var}(\widehat{\mathbf{Y}})^{-1}\operatorname{cov}(\widehat{\mathbf{Y}},Y)=\frac{N-n}{N}(1,0,\dots,0)^{\mathrm{\scriptscriptstyle T}}.$
Thus, we have $\widehat{\theta}^{\text{sada}}=N^{-1}\sum_{i=1}^{N}\widehat{y}_{1,i}$, and
$\mathbb{E}\{(\widehat{\theta}^{\text{sada}}-\theta^{*})^{2}\}=\frac{1}{N}\operatorname{var}(Y).$
Proof of Proposition 2..
The EIF for the outcome mean is a special case of Theorem 2(ii), whose proof is given in Supplementary Section S5, so we omit it here.
Assume $\widehat{Y}_{k}=\widehat{f}_{k}(\mathbf{X})$ for $k=1,\dots,K$, and $\widehat{Y}_{1}\equiv\mathbb{E}(Y\mid\mathbf{X})$. Let $\widehat{\mathbf{f}}_{-1}(\mathbf{x})=(\widehat{f}_{2}(\mathbf{x}),\dots,\widehat{f}_{K}(\mathbf{x}))^{\mathrm{\scriptscriptstyle T}}$. Then, we have
$\operatorname{var}(\widehat{\mathbf{Y}})=\left(\begin{array}[]{cc}\operatorname{var}\{\mathbb{E}(Y\mid\mathbf{X})\}&\operatorname{cov}\{\widehat{\mathbf{f}}_{-1}(\mathbf{X}),\mathbb{E}(Y\mid\mathbf{X})\}^{\mathrm{\scriptscriptstyle T}}\\ \operatorname{cov}\{\widehat{\mathbf{f}}_{-1}(\mathbf{X}),\mathbb{E}(Y\mid\mathbf{X})\}&\operatorname{var}\{\widehat{\mathbf{f}}_{-1}(\mathbf{X})\}\end{array}\right)=:\left(\begin{array}[]{cc}D_{11}&\bm{D}_{21}^{\mathrm{\scriptscriptstyle T}}\\ \bm{D}_{21}&\bm{D}_{22}\end{array}\right),$
and
$\operatorname{cov}(\widehat{\mathbf{Y}},Y)=\operatorname{cov}\left(\left(\begin{aligned} \mathbb{E}(Y\mid\mathbf{X})\\ \widehat{\mathbf{f}}_{-1}(\mathbf{X})\end{aligned}\right),Y\right)=\operatorname{cov}\left(\left(\begin{aligned} \mathbb{E}(Y\mid\mathbf{X})\\ \widehat{\mathbf{f}}_{-1}(\mathbf{X})\end{aligned}\right),\mathbb{E}(Y\mid\mathbf{X})\right)=:\left(\begin{aligned} D_{11}\\ \bm{D}_{21}\end{aligned}\right).$
Similar to (S2), by the inversion formula of $2\times 2$ block matrix, we obtain
$\bm{\omega}^{\text{opt}}=\frac{N-n}{N}\operatorname{var}(\widehat{\mathbf{Y}})^{-1}\operatorname{cov}(\widehat{\mathbf{Y}},Y)=\frac{N-n}{N}(1,0,\dots,0)^{\mathrm{\scriptscriptstyle T}}.$
Thus, we have $\widehat{\theta}^{\text{sada}}=n^{-1}\sum_{i=1}^{n}\{y_{i}-\mathbb{E}(Y\mid\mathbf{x}_{i})\}+N^{-1}\sum_{i=1}^{N}\mathbb{E}(Y\mid\mathbf{x}_{i})$, and
$\sqrt{N}(\widehat{\theta}^{\text{sada}}-\theta^{*})$ $=\frac{\sqrt{N}}{n}\sum_{i=1}^{n}\{y_{i}-\mathbb{E}(Y\mid\mathbf{x}_{i})\}+\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\mathbb{E}(Y\mid\mathbf{x}_{i})\}-\theta^{*}$
$=\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\phi_{\text{eif}}(r_{i},\mathbf{x}_{i},y_{i})+o_{p}(1).$
where $n/N\to\pi$ and
$\phi_{\text{eif}}(r,\mathbf{x},y)=\frac{r}{\pi}\{y-\mathbb{E}(Y\mid\mathbf{x})\}-\mathbb{E}(Y\mid\mathbf{x})-\theta^{*}.$
Appendix S5 Proof of Theorem 1
Proof.
Following arguments similar to those in the proof of Proposition 1 in Appendix S3, under assumption 1, we have
$\widehat{\bm{\theta}}^{\text{sada}}-{\bm{\theta}}^{*}$ (S3)
$~\doteq-\mathbf{H}^{-1}\left[\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}}^{*})+(\widehat{\mathcal{W}}^{\text{opt}})^{\mathrm{\scriptscriptstyle T}}\left\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})\right\}\right]$
$~=-\mathbf{H}^{-1}\left[\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}}^{*})+(\mathcal{W}^{\text{opt}})^{\mathrm{\scriptscriptstyle T}}\left\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})\right\}\right]$
$\qquad-\mathbf{H}^{-1}(\widehat{\mathcal{W}}^{\text{opt}}-\mathcal{W}^{\text{opt}})^{\mathrm{\scriptscriptstyle T}}\left\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})\right\}$
$~=:T_{1}+T_{2},$
where $\mathbf{H}=\mathbb{E}\{\partial\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}/\partial{\bm{\theta}}\}$. We next show that $T_{2}=o_{p}(n^{-1/2})$. Recall that
$\mathcal{W}^{\text{opt}}=\frac{N-n}{N}\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}^{-1}\mathbb{E}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}=:\frac{N-n}{N}\overline{\mathcal{W}}.$
Then,
$\|T_{2}\|\leq$ $~\left\|\mathbf{H}^{-1}\right\|\times\frac{N-n}{N}\left\|\widehat{\overline{\mathcal{W}}}-\overline{\mathcal{W}}\right\|\times\left\|\frac{1}{N-n}\sum_{i=n+1}^{N}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})\right\|$
$=$ $~O(1)\times o_{p}\left(\frac{N-n}{N}\right)\times O_{p}\left(\sqrt{\frac{1}{N-n}+\frac{1}{n}}\right)=o_{p}(n^{-1/2}),$
where the first equality is by $\widehat{\mathcal{W}}^{\text{opt}}\xrightarrow{p}\mathcal{W}^{\text{opt}}$, Chebyshev’s inequality and
$\mathbb{E}\left\|\frac{1}{N-n}\sum_{i=n+1}^{N}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})-\frac{1}{n}\sum_{i=1}^{n}\mathcal{S}(\mathbf{x}_{i},\widehat{\mathbf{y}}_{i};{\bm{\theta}}^{*})\right\|^{2}=\left(\frac{1}{N-n}+\frac{1}{n}\right)\left\|\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}\right\|.$
Moreover,
$\operatorname{var}(T_{1})=$ $~\mathbf{H}^{-1}\Bigg[\frac{1}{n}\operatorname{var}\{\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}}^{*})\}+\left(\frac{1}{N-n}+\frac{1}{n}\right)(\mathcal{W}^{\text{opt}})^{\mathrm{\scriptscriptstyle T}}\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}\mathcal{W}^{\text{opt}}$
$\qquad\qquad-\frac{2}{n}\operatorname{cov}\{\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}}^{*}),\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}\mathcal{W}^{\text{opt}}\Bigg]\mathbf{H}^{-1}$
$=$ $~\mathbf{H}^{-1}\left\{\frac{1}{n}{\bm{\Sigma}}_{\text{nv}}-\frac{N-n}{Nn}{\bm{\Sigma}}_{g}\right\}\mathbf{H}^{-1},$
where ${\bm{\Sigma}}_{\text{nv}}=\operatorname{var}\{\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}}^{*})\}$ and
${\bm{\Sigma}}_{g}=\mathbb{E}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}^{-1}\mathbb{E}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}.$
By the central limit theorem and the Slutsky’s theorem, we have
$\sqrt{n}(\widehat{\bm{\theta}}^{\text{sada}}-{\bm{\theta}}^{*})\xrightarrow{d}N\Big(\mathbf{0},\mathbf{H}^{-1}\big\{{\bm{\Sigma}}_{\text{nv}}-\frac{N-n}{N}{\bm{\Sigma}}_{g}\big\}\mathbf{H}^{-1}\Big).$
Appendix S6 Proof of Theorem 2
Proof.
Part (i). Without loss of generality, we assume $\widehat{Y}_{1}\equiv Y$. We denote $\mathbf{s}\equiv\mathbf{s}_{1}(\mathbf{X},\widehat{Y}_{1};{\bm{\theta}}^{*})\equiv\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})$ and $\mathbf{s}_{-1}\equiv(\mathbf{s}(\mathbf{X},\widehat{Y}_{2};{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}},\cdots,\mathbf{s}(\mathbf{X},\widehat{Y}_{K};{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. Then
$\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}=\left(\begin{array}[]{cc}\operatorname{var}(\mathbf{s}_{1})&\operatorname{cov}(\mathbf{s}_{-1},\mathbf{s}_{1})^{\mathrm{\scriptscriptstyle T}}\\ \operatorname{cov}(\mathbf{s}_{-1},\mathbf{s}_{1})&\operatorname{var}(\mathbf{s}_{-1})\end{array}\right),$
and $\mathbb{E}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}=(\operatorname{var}(\mathbf{s}_{1}),\operatorname{cov}(\mathbf{s}_{-1},\mathbf{s}_{1})^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. Similar to the proof in Appendix S4, by the inversion formula of block matrix, we have $\mathcal{W}^{\text{opt}}=(\mathbf{I},\mathbf{0},\dots,\mathbf{0})^{\mathrm{\scriptscriptstyle T}}\cdot(N-n)/N$, and
${\bm{\Sigma}}_{g}=$ $~\mathbb{E}\{\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}^{-1}\mathbb{E}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}$
$=$ $~\operatorname{var}(\mathbf{s}_{1})={\bm{\Sigma}}_{\text{nv}}.$
Then by Theorem 2, we have
$\widehat{\bm{\theta}}^{\text{sada}}-{\bm{\theta}}^{*}\xrightarrow{d}\mathcal{N}\left(\mathbf{0},\frac{1}{N}\mathbf{H}^{-1}{\bm{\Sigma}}_{\text{nv}}\mathbf{H}^{-1}\right).$
It is asymptotically equivalent to the oracle estimator who knows the ground truth of the unlabeled data and solves
$\frac{1}{N}\sum_{i=1}^{N}\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}})=0.$
Part (ii). We first derive the EIF for estimating $\bm{\theta}$$^{*}$, based on labeled data $\mathcal{L}=\{(\mathbf{x}_{i},y_{i}),i=1,\dots,n\}$ and unlabeled data $\mathcal{U}=\{\mathbf{x}_{i},i=n+1,\dots,N\}$. The joint density from one observation is
$\{\pi f(\mathbf{x})f(y\mid\mathbf{x})\}^{r}\{(1-\pi)f(\mathbf{x})\}^{1-r}=\pi^{r}(1-\pi)^{1-r}f(\mathbf{x})f(y\mid\mathbf{x})^{r}.$
It’s straightforward to show the tangent space of this model is $\mathcal{T}=\Lambda_{1}\bigoplus\Lambda_{2}$, where
$\Lambda_{1}=\left\{r\mathbf{b}(y,\mathbf{x}):\mathbb{E}(\mathbf{b}\mid\mathbf{x})=0\right\},~~\text{and}~~\Lambda_{2}=\left\{\mathbf{a}(\mathbf{x}):\mathbb{E}(\mathbf{a})=0\right\}.$
Let
$\bm{\Phi}_{\text{eif}}(r,\mathbf{x},y)=-\mathbf{H}^{-1}\left[\frac{r}{\pi}\{\mathbf{s}(\mathbf{x},y;{\bm{\theta}}^{*})-\bm{\mu}(\mathbf{x})\}+\bm{\mu}(\mathbf{x})\right],$
where $\bm{\mu}(\mathbf{x})=\mathbb{E}\{\mathbf{s}(\mathbf{x},Y;{\bm{\theta}}^{*})\mid\mathbf{x}\}$, $\frac{r}{\pi}\{\mathbf{s}(\mathbf{x},y;{\bm{\theta}}^{*})-\bm{\mu}(\mathbf{x})\}\in\Lambda_{1}$ and $\bm{\mu}(\mathbf{x})\in\Lambda_{2}$. It is easy to check $\bm{\Phi}_{\text{eif}}$ satisfies the orthogonal conditions (Tsiatis 2006, Theorems 4.2 and 4.3) and is thus the efficient influence function.
Without loss of generality, we assume $\mathbf{s}(\mathbf{x},\widehat{y}_{1};{\bm{\theta}}^{*})=\mathbf{s}(\mathbf{x},\widehat{f}_{1}(\mathbf{x});{\bm{\theta}}^{*})\equiv\bm{\mu}(\mathbf{x})$. We denote $\mathbf{s}_{-1}(\mathbf{x})=(\mathbf{s}(\mathbf{x},\widehat{f}_{2}(\mathbf{x});{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}},\dots,\mathbf{s}(\mathbf{x},\widehat{f}_{K}(\mathbf{x});{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. Then
$\operatorname{var}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\}=\left(\begin{array}[]{cc}\operatorname{var}\{\bm{\mu}(\mathbf{X})\}&\operatorname{cov}\{\mathbf{s}_{-1}(\mathbf{X}),\bm{\mu}(\mathbf{X})\}^{\mathrm{\scriptscriptstyle T}}\\ \operatorname{cov}\{\mathbf{s}_{-1}(\mathbf{X}),\bm{\mu}(\mathbf{X})\}&\operatorname{var}\{\mathbf{s}_{-1}(\mathbf{X})\}\end{array}\right),$
and $\mathbb{E}\{\mathcal{S}(\mathbf{X},\widehat{\mathbf{Y}};{\bm{\theta}}^{*})\mathbf{s}(\mathbf{X},Y;{\bm{\theta}}^{*})^{\mathrm{\scriptscriptstyle T}}\}=(\operatorname{var}\{\bm{\mu}(\mathbf{X})\},\operatorname{cov}\{\mathbf{s}_{-1}(\mathbf{X}),\bm{\mu}(\mathbf{X})\}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. Similar to the proof in Appendix S4, by the inversion formula of block matrix,, we have $\mathcal{W}^{\text{opt}}=(\mathbf{I},\mathbf{0},\dots,\mathbf{0})^{\mathrm{\scriptscriptstyle T}}\cdot(N-n)/N$. Then, by (S3), we have
$\sqrt{N}$ $\{\widehat{\bm{\theta}}^{\text{sada}}-{\bm{\theta}}^{*}\}$
$=$ $-\mathbf{H}^{-1}\left[\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}}^{*})+\frac{N-n}{N}\left\{\frac{1}{N-n}\sum_{i=n+1}^{N}\mathbf{s}(\mathbf{x}_{i},\widehat{y}_{1,i};{\bm{\theta}}^{*})-\frac{1}{n}\sum_{i=1}^{n}\mathbf{s}(\mathbf{x}_{i},\widehat{y}_{1,i};{\bm{\theta}}^{*})\right\}\right]$
$=$ $-\mathbf{H}^{-1}\left[\frac{\sqrt{N}}{n}\sum_{i=1}^{n}\{\mathbf{s}(\mathbf{x}_{i},y_{i};{\bm{\theta}}^{*})-\bm{\mu}(\mathbf{x}_{i})\}+\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\widehat{\mu}(\mathbf{x}_{i})\right]+o_{p}(1)$
$=$ $\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\bm{\Phi}_{\text{eif}}(r_{i},\mathbf{x}_{i},y_{i})+o_{p}(1),$
due to $n/N\to\pi$.
Appendix S7 Proof of Theorem 3
Proof.
Theorem 3 follows directly from theorem S4, which provides a non-asymptotic bound for the expected loss. ∎
Theorem S4.
(i) Under Assumption 2, we have that with probability at least $1-\delta$,
$\mathcal{R}(\widehat{{\bm{\theta}}}^{\text{nv}})-\mathcal{R}({\bm{\theta}}^{*})\leq\frac{16\sqrt{2}C}{\sqrt{n}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{1})}d\epsilon+\sqrt{\frac{24B^{2}}{n}\log\frac{2}{\delta}}+\frac{2B}{3n}\log\frac{2}{\delta},$
and
$\mathcal{R}(\widehat{\bm{\theta}}^{\text{sada}})$- $\mathcal{R}({\bm{\theta}}^{*})\leq\sqrt{\frac{128BC\sqrt{2n}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{3})}d\epsilon+8n\gamma B^{2}}{n^{2}}\log\frac{4}{\delta}}$
$+\frac{16\sqrt{2}C}{\sqrt{n}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{3})}d\epsilon+\frac{16\sqrt{2(N-n)}C}{N}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{2})}d\epsilon$
$+\sqrt{\frac{24B^{2}(N-n)}{N^{2}}\log\frac{4}{\delta}}+\frac{4B}{3N}\log\frac{4}{\delta},$
where $\gamma=\sup_{{\bm{\theta}}\in\Theta}\operatorname{var}(\widetilde{\ell}_{\bm{\theta}})/\operatorname{var}(\ell_{\bm{\theta}})\leq 1$, and $B$ is the upper bound of the loss function $\ell(f_{\bm{\theta}}(\mathbf{x}),y)$.
(ii) If additionally $\widehat{Y}_{k}=Y$ for some $k\in\{1,\dots,K\}$, we have that with probability at least $1-\delta$,
$\mathcal{R}(\widehat{\bm{\theta}}^{\text{sada}})$- $\mathcal{R}({\bm{\theta}}^{*})\leq\frac{16\sqrt{2}C}{\sqrt{N}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{1})}d\epsilon+\sqrt{\frac{24B^{2}}{N}\log\frac{2}{\delta}}+\frac{2B}{3N}\log\frac{2}{\delta},$
Proof of Theorem S4.
Recall that the simplified notations
$\ell_{\bm{\theta}}:=\ell(f_{\bm{\theta}}(\mathbf{x}),\mathbf{y}),~~\widehat{\ell}_{\bm{\theta}}:=\frac{N}{N-n}\bm{\omega}^{\mathrm{\scriptscriptstyle T}}\mathcal{L}(f_{\bm{\theta}}(\mathbf{x}),\widehat{\mathbf{y}}),~~\widetilde{\ell}_{\bm{\theta}}:=\ell_{\bm{\theta}}-\frac{N-n}{N}\widehat{\ell}_{\bm{\theta}},$
$\ell_{{\bm{\theta}},i}:=\ell(f_{\bm{\theta}}(\mathbf{x}_{i}),\mathbf{y}_{i}),~~\widehat{\ell}_{{\bm{\theta}},i}:=\frac{N}{N-n}\bm{\omega}^{\mathrm{\scriptscriptstyle T}}\mathcal{L}(f_{\bm{\theta}}(\mathbf{x}_{i}),\widehat{\mathbf{y}}_{i}),~~\widetilde{\ell}_{{\bm{\theta}},i}:=\ell_{{\bm{\theta}},i}-\frac{N-n}{N}\widehat{\ell}_{{\bm{\theta}},i}$
We can rewrite the loss functions as
$\widehat{\bm{\theta}}^{\text{nv}}=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\ell_{{\bm{\theta}},i},$
and
$\widehat{\bm{\theta}}^{\text{sada}}=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\widehat{\mathcal{R}}^{\text{sada}}({\bm{\theta}})=\operatornamewithlimits{arg\,min}_{{\bm{\theta}}\in\Theta}\frac{1}{n}\sum_{i=1}^{n}\widetilde{\ell}_{{\bm{\theta}},i}+\frac{N-n}{N}\frac{1}{N-n}\sum_{i=n+1}^{N}\widehat{\ell}_{{\bm{\theta}},i}.$
Notice that
$\mathcal{R}(\widehat{\bm{\theta}}^{\text{nv}})-\mathcal{R}({\bm{\theta}}^{*})$ $=\mathcal{R}(\widehat{\bm{\theta}}^{\text{nv}})+\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}}^{*})-\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}}^{*})-\mathcal{R}({\bm{\theta}}^{*})$
$\leq\mathcal{R}(\widehat{\bm{\theta}}^{\text{nv}})+\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}}^{*})-\widehat{\mathcal{R}}^{\text{nv}}(\widehat{\bm{\theta}}^{\text{nv}})-\mathcal{R}({\bm{\theta}}^{*})$
$\leq\sup_{{\bm{\theta}}\in\Theta}\{\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})-\mathcal{R}({\bm{\theta}})\}+\sup_{{\bm{\theta}}\in\Theta}\{\mathcal{R}({\bm{\theta}})-\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})\}.$
Denote $g(S):=\sup_{{\bm{\theta}}\in\Theta}\left\{\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})-\mathcal{R}({\bm{\theta}})\right\}$. By the Assumption 2, we know that $\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})-\mathcal{R}({\bm{\theta}})$ is a sub-Gaussian process with respect to the metric $\frac{Cd_{1}}{\sqrt{n}}$. Applying the classical Dudley’s entropy integral bound, we have
$\mathbb{E}g(S)\leq 8\sqrt{2}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,Cd_{1}/\sqrt{n})}d\epsilon$ $=8\sqrt{2}\int_{0}^{\infty}\sqrt{\log N(\epsilon\sqrt{n}/C,\Theta,d_{1})}d\epsilon$
$=\frac{8\sqrt{2}C}{\sqrt{n}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{1})}d\epsilon.$
Also, by the Assumption 2 and the Talagrand’s concentration inequality (Sen 2018, Theorem 8.7), the following holds with probability at least $1-{\delta}/{2}$,
$g(S)\leq\mathbb{E}[g(S)]+\sqrt{\frac{6B^{2}}{n}\log\frac{2}{\delta}}+\frac{B}{3n}\log\frac{2}{\delta}.$
We can bound $\sup_{{\bm{\theta}}\in\Theta}\{\mathcal{R}({\bm{\theta}})-\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})\}$ in the same way. Combining them together, we know it holds with probability at least $1-\delta$,
$\mathcal{R}(\widehat{\bm{\theta}}^{\text{nv}})-\mathcal{R}({\bm{\theta}}^{*})\leq\frac{16\sqrt{2}C}{\sqrt{n}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{1})}d\epsilon+\sqrt{\frac{24B^{2}}{n}\log\frac{2}{\delta}}+\frac{2B}{3n}\log\frac{2}{\delta}.$
We will decompose the SADA estimator $\widehat{\mathcal{R}}^{\text{sada}}({\bm{\theta}})$ into two terms and handle each of them using the same method as above,
$\widehat{\mathcal{R}}^{\text{sada}}({\bm{\theta}})=\frac{1}{n}\sum_{i=1}^{n}\widetilde{\ell}_{{\bm{\theta}},i}+\frac{N-n}{N}\frac{1}{N-n}\sum_{i=n+1}^{N}\widehat{\ell}_{{\bm{\theta}},i}=\widehat{\mathcal{R}}_{3}({\bm{\theta}})+\frac{N-n}{N}\widehat{\mathcal{R}}_{2}({\bm{\theta}}).$
where $\widehat{\mathcal{R}}_{2}({\bm{\theta}})={(N-n)}^{-1}\sum_{i=n+1}^{N}\widehat{\ell}_{{\bm{\theta}},i}$ and $\widehat{\mathcal{R}}_{3}({\bm{\theta}})={n}^{-1}\sum_{i=1}^{n}\widetilde{\ell}_{{\bm{\theta}},i}$. By the same argument as before,
$\mathcal{R}(\widehat{\bm{\theta}}^{\text{sada}})-\mathcal{R}({\bm{\theta}}^{*})$ $\leq\sup_{{\bm{\theta}}\in\Theta}\left\{\widehat{\mathcal{R}}^{\text{sada}}({\bm{\theta}})-\mathcal{R}({\bm{\theta}})\right\}+\sup_{{\bm{\theta}}\in\Theta}\left\{\mathcal{R}({\bm{\theta}})-\widehat{\mathcal{R}}^{\text{sada}}({\bm{\theta}})\right\}$
$\leq\sup_{{\bm{\theta}}\in\Theta}\left\{\widehat{\mathcal{R}}_{3}({\bm{\theta}})-\mathbb{E}[\widetilde{\ell}_{\bm{\theta}}]\right\}+\sup_{{\bm{\theta}}\in\Theta}\left\{\mathbb{E}[\widetilde{\ell}_{\bm{\theta}}]-\widehat{\mathcal{R}}_{3}({\bm{\theta}})\right\}$
$\quad+\frac{N-n}{N}\sup_{{\bm{\theta}}\in\Theta}\left\{\widehat{\mathcal{R}}_{2}({\bm{\theta}})-\mathbb{E}[\widehat{\ell}_{\bm{\theta}}]\right\}+\frac{N-n}{N}\sup_{{\bm{\theta}}\in\Theta}\left\{\mathbb{E}[\widehat{\ell}_{\bm{\theta}}]-\widehat{\mathcal{R}}_{2}({\bm{\theta}})\right\}.$
It follows from the exact same argument as $\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})$ that the following holds with probability at least $1-{\delta}/{2}$,
$\sup_{{\bm{\theta}}\in\Theta}\left\{\widehat{\mathcal{R}}_{2}({\bm{\theta}})-\mathbb{E}[\widehat{\ell}_{\bm{\theta}}]\right\}+\sup_{{\bm{\theta}}\in\Theta}\left\{\mathbb{E}[\widehat{\ell}_{\bm{\theta}}]-\widehat{\mathcal{R}}_{2}({\bm{\theta}})\right\}$ $\leq\frac{16\sqrt{2}C}{\sqrt{N-n}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{2})}d\epsilon$
$\quad+\sqrt{\frac{24B^{2}}{N-n}\log\frac{4}{\delta}}+\frac{2B}{3(N-n)}\log\frac{4}{\delta}.$
Denote $f(S):=\sup_{{\bm{\theta}}\in\Theta}\left\{\widehat{\mathcal{R}}_{3}({\bm{\theta}})-\mathbb{E}[\widetilde{\ell}_{\bm{\theta}}]\right\}$. Again, by the Dudley’s entropy integral bound, we have
$\mathbb{E}f(S)\leq 8\sqrt{2}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,Cd_{3}/\sqrt{n})}d\epsilon\leq\frac{8\sqrt{2}C}{\sqrt{n}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{3})}d\epsilon$
By the Talagrand’s inequality, it holds with probability $1-{\delta}/{4}$,
$f(S)\leq\mathbb{E}f(S)+\sqrt{\frac{32BC\sqrt{2n}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{3})}d\epsilon+2n\gamma B^{2}}{n^{2}}\log\frac{4}{\delta}}+\frac{B}{3n}\log\frac{4}{\delta}.$
We can bound $\sup_{{\bm{\theta}}\in\Theta}\left\{\mathbb{E}[\widetilde{\ell}_{\bm{\theta}}]-\widehat{\mathcal{R}}_{3}({\bm{\theta}})\right\}$ similarly. Putting all pieces together, it holds with probability at least $1-\delta$,
$\mathcal{R}(\widehat{\bm{\theta}}^{\text{sada}})-\mathcal{R}({\bm{\theta}}^{*})$ $\leq\sqrt{\frac{128BC\sqrt{2n}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{3})}d\epsilon+8n\gamma B^{2}}{n^{2}}\log\frac{4}{\delta}}$
$\quad+\frac{16\sqrt{2}C}{\sqrt{n}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{3})}d\epsilon+\frac{16\sqrt{2(N-n)}C}{N}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{2})}d\epsilon$
$\quad+\sqrt{\frac{24B^{2}(N-n)}{N^{2}}\log\frac{4}{\delta}}+\frac{4B}{3N}\log\frac{4}{\delta}.$
When there exists $k\in\{1,\dots,K\}$ so that $Y=\widehat{Y}_{k}$, we have $\widehat{\mathcal{R}}^{\text{sada}}({\bm{\theta}})=\frac{1}{N}\sum_{i=1}^{N}\ell_{{\bm{\theta}},i}$, which essentially has the same form as $\widehat{\mathcal{R}}^{\text{nv}}({\bm{\theta}})$ but replace $n$ with $N$. So following the same logic as before, we have
$\mathcal{R}(\widehat{\bm{\theta}}^{\text{sada}})$- $\mathcal{R}({\bm{\theta}}^{*})\leq\frac{16\sqrt{2}C}{\sqrt{N}}\int_{0}^{\infty}\sqrt{\log N(\epsilon,\Theta,d_{1})}d\epsilon+\sqrt{\frac{24B^{2}}{N}\log\frac{2}{\delta}}+\frac{2B}{3N}\log\frac{2}{\delta}.$
Appendix S8 Additional simulation results
Table S3 and S4 present the bias and coverage probability of the methods examined in the simulation studies described in Section 5.
Table S3: Bias of different methods under varying prediction quality $\gamma$.
Methods Bias
$\gamma=0$ $\gamma=0.1$ $\gamma=0.2$ $\gamma=0.3$ $\gamma=0.4$ $\gamma=0.5$ $\gamma=0.6$ $\gamma=0.7$ $\gamma=0.8$ $\gamma=0.9$ $\gamma=1$
Naive - 0.007 0.007 -0.001 -0.002 -0.001 0.006 -0.004 -0.009 0.002 0.009 -0.003
PPI $\widehat{Y}_{1}$ 0.005 0.009 -0.005 -0.002 -0.001 0.004 0.001 -0.004 0.000 0.002 0.003
$\widehat{Y}_{2}$ -0.001 0.000 -0.003 0.001 -0.003 0.000 -0.002 -0.001 0.005 0.010 0.001
PPI++ $\widehat{Y}_{1}$ 0.007 0.007 -0.001 -0.003 -0.002 0.005 0.000 -0.006 0.001 0.003 0.000
$\widehat{Y}_{2}$ 0.002 0.003 -0.003 0.000 -0.003 0.003 -0.003 -0.006 0.003 0.009 -0.003
SADA both 0.002 0.002 -0.003 -0.000 -0.003 0.003 0.000 -0.005 0.001 0.003 0.000
Table S4: Coverage probability of different methods under varying prediction quality $\gamma$.
Methods Relative efficiency
$\gamma=0$ $\gamma=0.1$ $\gamma=0.2$ $\gamma=0.3$ $\gamma=0.4$ $\gamma=0.5$ $\gamma=0.6$ $\gamma=0.7$ $\gamma=0.8$ $\gamma=0.9$ $\gamma=1$
Naive - .947 .940 .957 .951 .943 .930 .931 .953 .945 .955 .947
PPI $\widehat{Y}_{1}$ .948 .954 .957 .949 .939 .952 .947 .938 .950 .946 .935
$\widehat{Y}_{2}$ .962 .953 .942 .953 .946 .937 .939 .959 .951 .948 .954
PPI++ $\widehat{Y}_{1}$ .942 .939 .955 .954 .931 .935 .936 .941 .942 .958 .935
$\widehat{Y}_{2}$ .955 .944 .943 .945 .944 .929 .924 .950 .938 .951 .945
SADA both .946 .932 .939 .949 .938 .932 .938 .943 .937 .957 .924
Appendix S9 Generating predictions using LLMs
In this section, we provide a step-by-step description of how we generate predictions using the large language models like GPT-4o, Llama-3-8B, and DeepSeek in the empirical study presented in Section 6.1.
The input provided to the LLM consists of four components: a detail section that gives natural language descriptions of each data column, a background section that provides shared task context for all observations for example what kind of text is being evaluated and how politeness is defined, a question section that specifies what the model is asked to predict for every data point and describes the required output format for example a politeness score on a fixed scale., and a set of ten prompts, each designed to represent a different tone or inquiry style. As a result, each data point produces ten separate outputs, one for each prompt, and the final prediction is obtained by averaging these outputs.
Input prompt for generating responses Details: The text under review is: {text}, and is from the community: {source}. Please evaluate its level of politeness based on linguistic features. Background: {backgroud}. Question: {question}. Prompts: {prompt 1},…, {prompt 10}.
Below is an example of the prompts used in the politeness evaluation experiment.
An example prompt used in the politeness evaluation experiment Details: The text under review is: “Doing a redundant ‘atoi’ after you just did ‘strtol’ is probably the most twisted insult to the proper use of string-to-int conversion functions in C one can come up with. Why do you see the need to “re-convert” the value using the broken function ‘atoi’ when you already have it as ‘val’ from a proper function ‘strtol’?”, and is from the community: Stack Overflow. Please evaluate its level of politeness based on linguistic features. Background: This is a post from an online blog. Based on its linguistic features—such as word choice and tone—please evaluate the level of politeness. Then, assign a score from 1 (very impolite) to 25 (extremely polite), just like a sommelier would. Question: Choose an integer between 1 and 25 that reflects your evaluation. No explanation is needed, and do not give any numbers other than the rating. Your answer must be in JSON format with an integer only, without additional text. Prompts: Given the writing style and overall tone of this blog post, how would you rate its level of politeness? Taking into account the language used, the word choices, and the sentiment conveyed, what politeness score would you assign? Imagine you’re an expert in online communication. Based on the phrasing and attitude in this post, what number would you give for politeness? As someone who understands tone and subtext, how polite does this blog post feel to you? Just like a sommelier tasting wine, how would you score the politeness of this post, based solely on its language? If you had to assign a politeness score between 1 and 25 to this writing sample, what would it be? What’s your judgment on the tone of this blog post—how polite does it seem on a scale from 1 to 25? Considering the linguistic subtleties in this post—formality, tone, and word choice—how would you evaluate its politeness? Reflecting on how this post might come across to a casual reader, what politeness score would you give it? Evaluate the post as if you were rating its etiquette. What score feels most appropriate?